### v1

# Revealing Language Model Trajectories via Kullback-Leibler Divergence

Abstract
A recently proposed method enables efficient estimation of the KL divergence between language models, including models with different architectures, by assigning coordinates based on log-likelihood vectors. To better understand the behavior of this metric, we systematically evaluate KL divergence across a wide range of conditions using publicly available language models. Our analysis covers comparisons between pretraining checkpoints, fine-tuned and base models, and layers via the logit lens. We find that trajectories of language models, as measured by KL divergence, exhibit a spiral structure during pretraining and thread-like progressions across layers. Furthermore, we show that, in terms of diffusion exponents, model trajectories in the log-likelihood space are more constrained than those in weight space.
Revealing Language Model Trajectories via Kullback-Leibler Divergence
Ryo Kishino1  Yusuke Takase1  Momose Oyama1,2 Hiroaki Yamagiwa1   Hidetoshi Shimodaira1,2 1Kyoto University 2RIKEN kishino.ryo.32s@st.kyoto-u.ac.jp, {y.takase, oyama.momose}@sys.i.kyoto-u.ac.jp, {h.yamagiwa, shimo}@i.kyoto-u.ac.jp
1 Introduction
Figure 1: Visualization of the training trajectories of Pythia models using t-SNE applied to log-likelihood vectors. Each point represents a model checkpoint, with lighter colors indicating earlier steps. The thickness of the connecting lines reflects the magnitude of KL divergence between checkpoints. The top panel shows trajectories for five model sizes and multiple random seeds. The bottom panel shows only the five sizes, with both 2D and 3D t-SNE visualizations for clarity. See Section 3 for details.
between checkpoints between seeds between model sizes
10$\to$11k 50$\to$51k 100$\to$101k avg. 1B 1.4B 2.8B 6.9B
410M 0.069 0.053 0.023 0.11 0.28 0.52 1.0 1.7
($\pm$0.0011) ($\pm$0.0010) ($\pm$0.00042) ($\pm$ 0.0024) ($\pm$0.010) ($\pm$0.018) ($\pm$0.036) ($\pm$0.060)
(a) KL divergence for Pythia 410M, measured between checkpoints saved during pretraining, different random seeds, and different model sizes. For the inter-size comparison, the final checkpoints were used. See Section 3 for details.
Type (#pair) Pair Median Mean SD
mistral (25) fine-tuning 0.32 0.37 0.32
random 0.79 0.74 0.39
llama-3 (20) fine-tuning 0.20 0.67 0.80
random 0.66 0.87 0.66
all (66) fine-tuning 0.40 1.5 3.6
random 2.1 5.0 9.5
(b) Summary statistics (median, mean, SD) of KL divergence between fine-tuned and base models, and between random model pairs. See Section 4 for details.
Model 1$\to$2 16$\to$17 31$\to$32
Llama-2-7b-hf 0.61 1.6 1.3
(± 0.013) (± 0.063) (± 0.10)
Meta-Llama-3-8B 4.0 22 1.7
(± 0.075) (± 0.78) (± 0.037)
Mistral-7B-v0.3 2.4 5.7 8.6
(± 0.10) (± 0.12) (± 0.54)
(c) KL divergence between adjacent layers within models, measured using the logit lens. See Section 5 for details.
Table 1: KL divergence in bits per byte between language models under various conditions. Values in parentheses represent standard errors. See Appendix C for the formula used to compute the standard errors.
The growing diversity of language models has created a need to quantify their differences systematically. Treating them as probability distributions over text, the Kullback-Leibler (KL) divergence serves as a natural comparison metric. Oyama et al. (2025) proposed using log-likelihood vectors on a fixed text set to assign model coordinates in a Euclidean space. In this space, the squared distance between models approximates their KL divergence, making model comparison a geometric problem.
However, their analysis did not cover the typical KL divergence values. As an early report on model analysis leveraging log-likelihood vectors, we use publicly available models to measure the KL divergence between checkpoints saved during pretraining (Section 3), fine-tuned and base models (Section 4), and different layers (Section 5) as shown in Table 1.
Furthermore, we analyze the trajectories of language models throughout training and across layers, yielding several notable observations. As shown in Fig. 1, portions of the pretraining trajectories exhibit a spiral structure. Moreover, from the perspective of anomalous diffusion, model dynamics in the log-likelihood space become increasingly constrained during the later steps of training, compared to those in weight space.
2 Review: Model Map
Log-likelihood vector.
The probability that a language model $p$ generates a text (i.e., a sequence of tokens) $x=(y_{0},\ldots,y_{M})$ is given by $p(x)=\prod_{m=1}^{M}p(y_{m}\mid y_{0},\ldots,y_{m-1})$. Oyama et al. (2025) defined the log-likelihood vector of a language model $p$ over a predefined text set $\{x_{1},\ldots,x_{N}\}$ as
$\bm{\ell}=(\log p(x_{1}),\ldots,\log p(x_{N}))^{\top}\in\mathbb{R}^{N},$
and showed that it serves as a useful feature representation for comparing models and predicting downstream task performance.
KL divergence.
Let $\{p_{i}\}_{i=1}^{K}$ be a set of language models, and $\bm{L}=(\bm{\ell}_{1},\ldots,\bm{\ell}_{K})^{\top}\in\mathbb{R}^{K\times N}$ be the matrix formed by stacking the log-likelihood vectors $\bm{\ell}_{i}$ of each model. Define $\bm{Q}=(\bm{q}_{1},\ldots,\bm{q}_{K})^{\top}$ as the matrix obtained by double-centering [^1] Row-wise centering followed by column-wise centering. $\bm{L}$. Then, the KL divergence between two models $p_{i}$ and $p_{j}$ can be approximated as
$2\mathrm{KL}(p_{i},p_{j})\approx\|\bm{q}_{i}-\bm{q}_{j}\|^{2}/N.$
In other words, by associating each model $p_{i}$ with a coordinate $\bm{q}_{i}\in\mathbb{R}^{N}$, KL divergence is approximated by squared Euclidean distance in this space. This coordinate system is referred to as the model map in Oyama et al. (2025). In particular, we refer to a sequence of models represented as coordinates in the model map as a trajectory.
General experimental settings.
For the text set, we use the same 10,000 texts extracted from the Pile Gao et al. (2020) as in Oyama et al. (2025) [^2] https://github.com/shimo-lab/modelmap/. The values of KL divergence are normalized by the average text length, $\bar{B}=972.3188$ bytes, and are reported in units of bits per byte. In particular, the matrix $\bm{Q}$ is divided by $\sqrt{2N\bar{B}\log 2}$ in advance, so that the squared Euclidean distance directly approximates the KL divergence in bits per byte. All visualizations are based on this $\bm{Q}$.
3 Pretraining
In the Pythia suite Biderman et al. (2023), the checkpoints saved during pretraining are publicly available for each model size. Additionally, van der Wal et al. (2025) released checkpoints for smaller Pythia models trained with different random seeds. In this section, we investigate how a language model’s trajectory evolves during pretraining by analyzing these checkpoints.
Settings.
We use checkpoints from the standard Pythia models with sizes 410M, 1B, 1.4B, 2.8B, and 6.9B. In addition, for the 410M model, we supplement our analysis with checkpoints from seven different random seeds. The checkpoints are saved at training steps 0, 1, 2, 4, …, 512, 1k, 2k, 3k, …, and 143k. We excluded the 1B checkpoint at step 116k due to anomalous weights, and removed texts with irregular log-likelihood variations across checkpoints. See Appendix D.1 for details.
KL divergence.
Table 1(a) shows the KL divergence for Pythia 410M. The KL divergence between checkpoints saved every 1k steps gradually decreases as training progresses, and after step 10k, it stays within a small range (approximately 0.02 to 0.07 bit/byte), as also shown in the right panel of Fig. 4. Moreover, the values between different random seeds are around 0.1 bit/byte, while those between different sizes range from 0.3 to 2 bit/byte, indicating larger differences. See Appendix D.2 for details.
Trajectory.
Figure 2: Trajectory of Pythia 2.8B during pretraining, visualized by applying 3D t-SNE, followed by PCA on the later half portion. (Left) PC2 vs. PC3. (Right) Autocorrelation function of PC3.
Model Size 410M 1B 1.4B 2.8B 6.9B
Period (step) 30k 13k 15k 14k 12k
Table 2: Period of the spiral trajectory of Pythia during pretraining. See Appendix D.3 for details.
Fig. 1 shows a simultaneous visualization of the pretraining trajectories of five model sizes using t-SNE van der Maaten and Hinton (2008) [^3] We used the scikit-learn implementation with the default perplexity value of 30., which is not possible with weight-based representations. The trajectories evolve continuously during training [^4] The apparent jumps in the 2D trajectories of the 2.8B and 6.9B models are artifacts of t-SNE, which tends to separate overlapping points. These trajectories appear continuous in 3D t-SNE visualizations.. Moreover, the 3D visualization suggests that the trajectories exhibit a spiral structure. As shown in Fig. 2, applying PCA to the later portion of the trajectory obtained via t-SNE reveals a circular pattern. The autocorrelation function of PC3 exhibits periodicity. The estimated periods of the spiral trajectories are shown in Table 2. Similar phenomena in weight trajectories have been studied previously Cohen et al. (2021); Netay (2024) described in detail in Appendix A. In addition, refer to Appendix D.4 for a trajectory visualization using PCA.
Comparison with weights.
Figure 3: Temporal evolution of the squared Euclidean distance in the weights and log-likelihood vectors from step 120k to step 130k for Pythia. The diffusion exponent $c$ is estimated using the least squares method.
Figure 4: (Left) Diffusion exponent $c$ as a function of the starting point $t_{0}$. (Right) KL divergence between consecutive checkpoints of Pythia (corresponding to Table 1(a)). See Appendix D.5 for details.
Differences between models with the same architecture can be measured by the distance between their weights. Let $\bm{W}_{t}$ be the weights of a neural network at step $t$. According to Kunin et al. (2024), after model performance has converged, the weights diffuse following a power law with exponent $c$ [^5] Brownian motion corresponds to $c=1$, while anomalous diffusion is defined as $c\neq 1$. Specifically, $c>1$ means superdiffusion, and $c<1$ means subdiffusion. If the evolution is fractional Brownian motion Mandelbrot and Van Ness (1968), the Hurst exponent corresponds to $H=c/2$.. That is, with respect to an initial step $t_{0}$, the relationship $\|\bm{W}_{t}-\bm{W}_{t_{0}}\|^{2}\propto|t-t_{0}|^{c}$ holds, where $c$ may differ from $1$.
To verify this, we measured the evolution of the squared Euclidean distance in both the weights $\bm{W}$ and the log-likelihood vectors $\bm{q}$, over steps 120k to 130k of Pythia. As shown in Fig. 3, both $\bm{W}$ and $\bm{q}$ follow a power-law relationship with exponents $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ respectively. The weights $\bm{W}$ exhibit Brownian-like diffusion ($c_{\mathrm{w}}\approx 1$), whereas the diffusion of $\bm{q}$ is substantially suppressed ($c_{\mathrm{q}}\approx 0.2$). In addition, Fig. 4 illustrates how the exponents $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ vary with the starting step. Notably, $c_{\mathrm{q}}$ remains consistently smaller than $c_{\mathrm{w}}$ throughout. Both $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ tend to decrease as training progresses. However, around step 120k, when the KL divergence between successive checkpoints stabilizes, $c_{\mathrm{w}}$ increases, while $c_{\mathrm{q}}$ shows no significant change. As discussed in Appendix B, the difference in dynamics between these two spaces could be explained by the flat minima hypothesis Keskar et al. (2017) and by constraints imposed by the loss function on the log-likelihood space.
4 Fine-tuning
Figure 5: Map showing base models and their fine-tuned variants. Lines represent parent-child relationships between models. Colors indicate model types as defined in Oyama et al. (2025). Generation refers to the depth from the root model within each lineage.
The introduction of efficient fine-tuning methods Hu et al. (2022); Rafailov et al. (2023) has led to the development of a large number of fine-tuned models. In this section, we investigate the KL divergence between fine-tuned and base models.
Settings.
Out of the 1,018 language models analyzed in Oyama et al. (2025), we use 66 pairs of fine-tuned and base models (87 models in total; Fig. 5). The parent-child relationships are identified using Hugging Face Hub API. See Appendix E for details.
KL divergence.
Table 1(b) shows the statistics on the KL divergence between fine-tuned and base models, and between randomly selected models. For the cumulative distribution, see Fig. 18 in Appendix E. KL divergence values tend to be slightly smaller for models of type mistral or llama-3. In all cases, the KL divergence between fine-tuned and base models is smaller than that between random pairs, with a median ranging from 0.2 to 0.4 bit/byte.
5 Layer
Figure 6: t-SNE visualization of trajectories across layers for 50 models, obtained by applying the logit lens to each layer. Deeper layers are represented with lighter colors, and the final layer is indicated by a white square.
Figure 7: For 10 models per model type: (left) t-SNE visualization of trajectories across layers, where the final layer is indicated by a white square; (right) KL divergence between adjacent layers (corresponding to Table 1(c)). See Fig. 20 in Appendix F for a 3D plot.
We investigate model trajectories across layers by treating the network up to each layer as an individual model using the logit lens nostalgebraist (2020).
Settings.
We use a subset of the 1,018 language models analyzed in Oyama et al. (2025). First, without restricting model types, we select the 50 most downloaded models (Fig. 6). When limiting to specific model types, we use the 10 most downloaded models from each of the llama-2, llama-3, and mistral with 32 layers (30 models in total; Fig. 7). See Appendix F for details on the models used.
KL divergence.
The right panel of Fig. 7 and Table 1(c) show the KL divergence between consecutive layers. Models of the same type exhibit similar trends. For llama-2, Wendler et al. (2024) reported a sharp drop in the entropy of next-token distributions at intermediate layers in a translation task, which corresponds to the peak positions in Fig. 7. The KL divergence between consecutive layers ranges from as low as 1 bit/byte to as high as 10 to 25 bit/byte, indicating substantial variation.
Trajectory.
Fig. 6 and the left panel of Fig. 7 illustrate the model trajectories across layers. The trajectories exhibit jumps at KL divergence peaks and tend to align in deeper layers. The trajectories align with increasing mean log-likelihood (Fig. 19, Appendix F).
6 Conclusion
We analyzed KL divergence and trajectories across checkpoints, model pairs, and layers, revealing spiral and aligned structures over training and depth.
Limitations
  • The checkpoints for Pythia are available only at 1k-step intervals except for the early stages of training. When checkpoints are saved at shorter intervals, the resulting trajectories can differ from our observation. In particular, the estimation of the diffusion exponent may be affected by errors introduced by the coarse granularity of the 1k-step resolution.
  • The KL divergence we measured depends on the text set used. In future work, we plan to investigate how the KL divergence and model trajectories vary when a different text set is used.
  • The number of models used is limited in the experiment measuring KL divergence between fine-tuned and base models. This is because some fine-tuned models do not include information about their base models in their model cards obtained by using Hugging Face Hub API.
  • In the layer-wise analysis, the logits obtained from shallow layers via the logit lens contain substantial noise. This issue could be addressed by using the tuned lens Belrose et al. (2023), an improved version of the logit lens. However, the number of publicly available pretrained tuned lenses is currently limited. We believe this limitation can be overcome in the future by training tuned lenses ourselves.
  • While we observe structured behaviors such as spiral trajectories and subdiffusion in the log-likelihood space, their underlying causes and implications for learning dynamics have not been quantitatively analyzed and remain open for future investigation.
Acknowledgments
We thank Luis Iván Hernández Ruíz for helpful discussions. This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST BOOST JPMJBS2407.
Appendix A Related Work on Oscillatory Trajectories in Weight Space
We analyzed model trajectories in the log-likelihood space, where each model was represented by its output probabilities on a fixed text set. This space reflects the model’s behavior directly, and our study was, to the best of our knowledge, the first to reveal structural patterns in training dynamics within this representation. In particular, we observed the emergence of spiral-shaped trajectories as well as indications of constrained diffusion, which we further examine in Appendix B.
While our main focus was on the log-likelihood space, we also briefly examined model trajectories in weight space for comparison (Appendix D.5). As shown in Fig. 16, oscillatory behavior was also observed in the weight space during the later stages of training. Although the patterns differed in periodicity and structure, the presence of oscillations in both spaces suggests a potentially broader phenomenon in the dynamics of neural networks. The observed difference may reflect the presence of multiple modes of variation, with different modes being emphasized in the weight and output spaces.
To contextualize these findings, we review prior work that has reported oscillatory behaviors in weight space during training. The Edge of Stability phenomenon Cohen et al. (2021); Arora et al. (2022) refers to the behavior of training loss oscillating while decreasing, observed when gradient descent enters an unstable region where the largest Hessian eigenvalue hovers near the critical learning rate. Singh et al. (2025) observed micro-level oscillations in training trajectories through changes in the cosine similarity of weights. Netay (2024) empirically identified double-helical trajectories in neural network weights, noting a relationship between these trajectories’ periodicity and the hyperparameters of the Adam optimizer. Kunin et al. (2024) found that weights exhibit anomalous diffusion after convergence and demonstrated oscillatory dynamics in the weight phase space, governed by the eigenstructure of the Hessian of the loss function.
Appendix B Related Work and Discussions on Diffusion Exponents
In Section 3, we estimated the diffusion exponents $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ corresponding to the trajectories of models in weight and log-likelihood spaces, respectively. Our analysis revealed that during the early phase of training, both exponents are relatively large ($c_{\mathrm{w}}=1.2$ and $c_{\mathrm{q}}=0.8$), but they decrease as training progresses. At the later stages, the exponent in weight space approaches $c_{\mathrm{w}}\approx 1$, which corresponds to Brownian motion, while the exponent in log-likelihood space becomes significantly smaller ($c_{\mathrm{q}}\ll 1$), which is characteristic of subdiffusion.
In our experimental setting, texts were sampled from a corpus that closely approximates the training data. Consequently, each element of the log-likelihood vector represents a term contributing to the cross-entropy loss, and the negative average of these terms serves as an unbiased estimator of the expected loss. The contrast between the relatively large diffusion exponent in weight space and the much smaller exponent in log-likelihood space suggests that even as model weights continue to change, the output probabilities remain stable. As a result, the loss function is largely unaffected by these weight changes. This observation aligns with the hypothesis that stochastic gradient descent favors flat minima Keskar et al. (2017).
Moreover, near the loss minima, the sum of negative log-likelihoods is constrained as cross-entropy loss, implying convergence to regions where many texts have high log-likelihood values. The observed small diffusion exponent suggests that the movement of the vectors becomes restricted in these regions.
Appendix C Standard Error of the Estimated KL Divergence
Let $\bm{q}_{i}$ and $\bm{q}_{j}$ denote the $\bm{q}$ coordinates of models $p_{i}$ and $p_{j}$, respectively. While we use $\bm{q}$ pre-scaled by $1/\sqrt{N}$ in the main body of this paper, we do not apply this scaling in this appendix. Accordingly, the KL divergence between the models is estimated as $\mathrm{KL}(p_{i},p_{j})\approx\frac{1}{N}\|\bm{q}_{i}-\bm{q}_{j}\|^{2}=\frac{1}{N}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{2},$ which corresponds to the sample mean of $\{(Q_{ik}-Q_{jk})^{2}\}_{k=1}^{N}$. Assuming that the effects of double-centering in $\bm{Q}$ are negligible, these terms can be treated as approximately independent. When $N$ is sufficiently large, the variance of the estimated KL divergence is given by
$\mathrm{Var}(\mathrm{KL}(p_{i},p_{j}))$
$\approx\frac{1}{N}\left\{\frac{1}{N}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}-\frac{1}{N^{2}}\|\bm{q}_{i}-\bm{q}_{j}\|^{4}\right\}$
$\approx\frac{1}{N^{2}}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}-\frac{1}{N}\mathrm{KL}(p_{i},p_{j})^{2}.$
Therefore, the standard error of the estimator $\mathrm{KL}(p_{i},p_{j})$ can be approximated as
$\mathrm{SE}(\mathrm{KL}(p_{i},p_{j}))$
$\approx\sqrt{\frac{1}{N^{2}}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}-\frac{1}{N}\mathrm{KL}(p_{i},p_{j})^{2}}.$
Appendix D Details of Section 3
To analyze the trajectories of language models during pretraining, we used Pythia models [^6] Released under the Apache 2.0 License. of various sizes, as listed in Table LABEL:tab:model_list_pretraining in Appendix G. For each checkpoint, we computed log-likelihood vectors using the same set of 10,000 texts as in Oyama et al. (2025). The computations were performed on an NVIDIA RTX 6000 Ada and took approximately 10 minutes for a 7B model loaded in float16 precision.
D.1 Preprocessing
Figure 8: Squared Euclidean distance between the weights of consecutive checkpoints for Pythia 1B. The distances between 115k and 116k, and between 116k and 117k are abnormally large.
Figure 9: The relationship between the maximum and the standard deviation of the absolute differences in log-likelihoods between consecutive checkpoints of the same model size, computed for each text.
Figure 10: Change in KL divergence after step 10k as a function of the number of top outlier texts removed (10, 100, 200, …, 1000). Brighter colors indicate larger numbers of removed texts.
Assuming that the KL divergence between consecutive checkpoints changes smoothly, we regarded certain checkpoints and texts as outliers. All experiments were conducted after removing these outliers. This appendix provides a detailed description of the preprocessing procedure.
Removing outlier models.
As shown in Fig. 8, the Euclidean distance between weights near step 116k of Pythia 1B is abnormally large. We suspect that this checkpoint may not have been saved properly and therefore excluded it from our analysis. A question about this checkpoint was also raised on the Discussions page of the Pythia 1B model on Hugging Face [^7] https://huggingface.co/EleutherAI/pythia-1b/discussions/4.
Removing outlier texts.
Following Oyama et al. (2025), we clip the log-likelihood matrix $\bm{L}$ at the bottom 2%, after excluding the outlier model. Next, we consider the log-likelihood difference for each text between consecutive checkpoints after the warmup phase. For each text, we define its outlier score as the maximum absolute difference across all model sizes and training steps. Specifically, letting $p_{i,t}$ denote the model of size $i$ at step $t$, the outlier score for a text $x$ is given by $\max_{i,t}\left\{\left|p_{i,t+1}(x)-p_{i,t}(x)\right|\right\}$. It is also possible to assign scores to texts using the standard deviation of these differences. However, Fig. 9 shows that the maximum and the standard deviation are highly correlated. Therefore, we adopt the maximum value for scoring. The figure further confirms the presence of texts that clearly behave as outliers. For example, the text with the highest outlier score was a meaningless and repetitive sequence of symbols, such as 028a28a0028a28a0028a28a....
Fig. 10 shows the KL divergence between consecutive checkpoints after step 10k [^8] The maximum is computed from step 2k onward, after the warmup phase, but we visualize from step 10k to better capture the variation in KL divergence., plotted as a function of the number of top outlier-scored texts removed (top 10, 100, 200, …, 1000). When the top 300 texts (3%) are removed, the variation in KL divergence stabilizes with respect to the number of removed texts. Therefore, we excluded these 300 texts from our experiments.
D.2 KL Divergence between Checkpoints
(a) KL divergence between consecutive checkpoints during the warmup phase. The horizontal axis is scaled as $\log_{10}(1+\mathrm{Step})$.
(b) KL divergence between consecutive checkpoints after the warmup phase. When the vertical axis is plotted on a logarithmic scale, this corresponds to the right panel of Fig. 4.
Figure 11: KL divergence between consecutive checkpoints of Pythia. For example, the value at step 10k on the horizontal axis represents the KL divergence between the checkpoints at steps 10k and 11k.
Figure 12: KL divergence and its standard error between consecutive checkpoints of Pythia after step 10k.
410M 1B 1.4B 2.8B 6.9B
410M 0.00 0.28 0.52 1.04 1.7
1B 0.28 0.00 0.15 0.44 0.92
1.4B 0.52 0.15 0.00 0.21 0.57
2.8B 1.0 0.44 0.21 0.00 0.21
6.9B 1.7 0.92 0.57 0.21 0.00
Table 3: KL divergence (bit/byte) between different model sizes at the final checkpoint of Pythia.
Fig. 11 shows the variation in KL divergence between consecutive checkpoints for all model sizes. In this figure, the value at step 10k, for example, represents the KL divergence between the checkpoints at steps 10k and 11k. During pretraining, the learning rate changes dynamically with the training step. In particular, the first 1,430 steps correspond to the warmup phase, during which the learning rate increases linearly. As a result, as shown in Fig. 11, the KL divergence between the checkpoints takes on relatively large values, mostly exceeding 10 bit/byte during this interval. Additionally, Fig. 12 plots the KL divergence between consecutive checkpoints after step 10k for Pythia, along with the corresponding standard error. The derivation of the standard error is described in Appendix C. Finally, Table 3 presents the pairwise KL divergence between model sizes at the final checkpoint. As discussed in Section 3, the divergence caused by differences in model size is larger than that resulting from 1k steps of training.
D.3 Period of Spiral Trajectory
Figure 13: (Left) PC2 vs. PC3 after applying PCA to the later half portion of the 3D t-SNE trajectories for each Pythia model size. (Right) Autocorrelation function of PC3.
In this appendix, we describe the procedure used to analyze the spiral trajectories suggested in Fig. 1. For each model size in Pythia, 3D t-SNE is applied independently. PCA is then performed on the sequence of checkpoints from the midpoint of the trajectory (step 67k onward) to extract principal axes. Since PC1 corresponds to the overall direction of trajectory progression, we plot PC2 versus PC3 in the left panel of Fig. 13, where the trajectory approximately forms a circular pattern. We also compute the autocorrelation function of the PC3 values, treated as a time series, and show the result in the right panel of Fig. 13. The spiral period reported in Table 2 is defined as the step at which the autocorrelation reaches its maximum between the second and third zero-crossings.
D.4 PCA
Figure 14: PCA visualization of the training trajectories of Pythia models. (Top) PC1 vs. PC2. (Bottom) PC11 vs. PC12.
Figure 15: Variation of principal component values from PCA applied to the trajectories of all Pythia model sizes.
In Section 3, we visualized the trajectories of Pythia during pretraining using t-SNE. In this appendix, we instead visualize the trajectories using PCA.
Fig. 14 shows the result of applying PCA simultaneously to all model sizes and plotting their trajectories. Fig. 15 further illustrates the evolution of multiple principal component values along the training steps. Unlike t-SNE, the top principal components exhibit similar values across model sizes, resulting in overlapping trajectories. This difference likely arises from the fact that the top principal components capture global variation in the trajectories, whereas t-SNE emphasizes local structure. In contrast, the lower principal components exhibit more noise-like behavior. As shown in Table 3, the KL divergence between different model sizes at the final checkpoint is larger than that between consecutive checkpoints. This is consistent with the observation that trajectories are separated in the lower components.
D.5 Comparison with Weights
Figure 16: t-SNE visualization of the trajectories based on weight-space distances for Pythia models after step 100k. To align scales across model sizes, the trajectories were centered after t-SNE. They were also rotated to align the directions of progression. Final checkpoints are indicated using distinct markers.
Size Log-likelihood Weight
410M 0.849 0.993
($\pm$0.158) ($\pm$0.00622)
1B 0.935 0.999
($\pm$ 0.0866) ($\pm$0.000922)
1.4B 0.955 0.999
($\pm$ 0.0336) ($\pm$0.000687)
2.8B 0.959 0.999
($\pm$ 0.088) ($\pm$0.000704)
6.9B 0.977 0.999
($\pm$ 0.0305) ($\pm$0.000975)
Table 4: The mean and standard deviation of the coefficient of determination ($R^{2}$) when varying the starting step used to compute the diffusion exponent.
This appendix describes the procedure for comparing language models using the Euclidean distance between the weights of Pythia checkpoints.
Trajectory.
For each model size, we computed a distance matrix containing the pairwise Euclidean distances among all checkpoints. Fig. 16 shows the trajectories from step 100k onward, visualized using t-SNE999t-SNE can take a distance matrix as input.. The trajectories were centered after t-SNE to align the plot ranges. In addition, for the 1B model, the trajectory was rotated by 180 degrees to align its rough direction of progression with that of the other model sizes.
Anomalous diffusion.
Fig. 3 shows the evolution of squared Euclidean distances in weights and log-likelihood vectors over 10k steps starting from step 120k. Fig. 4 illustrates how the estimated diffusion exponent changes when varying the starting step for the calculation. To ensure stability, we vary the starting point only after the warmup phase. The exponent is estimated using linear regression (least squares) over a 10k-step window beginning at each initial step. Summary statistics of the coefficient of determination ($R^{2}$) for each regression are provided in Table 4.
D.6 Different Seeds
Figure 17: KL divergence between the final checkpoints of Pythia 410M models with different seeds.
van der Wal et al. (2025) trained Pythia models with nine different seeds and released checkpoints at the same steps as in the original Pythia suite. The random seed affects the initialization of weights and the order of the training data. We used all available checkpoints for the 410M models with seeds 1 through 9. However, at the time of writing this paper, the weights for (seed, step) = (2, 111k), (6, 88k), and (9, 40k) were not available on Hugging Face and were therefore excluded from our analysis [^10] For example, see https://huggingface.co/EleutherAI/pythia-410m-seed2/tree/step111000.. The original 410M model was trained with seed 1234.
Fig. 17 shows the KL divergence between the final checkpoints of models trained with different seeds. Despite sharing the same architecture and data, the KL divergence between seed 3 and the other seeds, as well as between seed 4 and the others, are significantly larger. According to van der Wal et al. (2025), models trained with seed 3 and seed 4 experienced loss spikes during training and are considered outliers in terms of performance. This suggests that KL divergence is also effective in identifying such outlier models. Therefore, seed 3 and seed 4 are excluded when generating Fig. 1. We also report in Table 1(a) the average KL divergence between the final checkpoint of the original 410M model and those of seeds 1, 2, 5, 6, 7, 8, and 9.
Appendix E Details of Section 4
Figure 18: Cumulative distribution of KL divergence for pairs of fine-tuned models and their base models, and for randomly selected model pairs (corresponding to Table 1(b)). The horizontal axis is scaled as $\log_{10}(1+\mathrm{KL})$. The dotted lines indicate the median of each distribution.
We describe the language models used in Section 4. From the 1,018 models analyzed in Oyama et al. (2025), we identified fine-tuned models as those whose base model is specified in the metadata available via the Hugging Face Hub API [^11] https://github.com/huggingface/huggingface_hub. Our analysis was conducted on pairs of fine-tuned models and their base models. Models associated with multiple base models (i.e., merged models) were excluded. A list of the language models used in this analysis is provided in Table LABEL:tab:model_list_ft in Appendix G. In Fig. 5, there are a few pairs connected by line segments that span across different colors, namely different model types. This occurs because, under the model type classification method used in Oyama et al. (2025), some fine-tuned models were categorized as belonging to the “others” type. In our experiments, when we refer to, for example, a “llama-3 fine-tuned pair”, we mean a pair in which the fine-tuned model is of the llama-3 type.
Fig. 18 illustrates the cumulative distributions of KL divergence for pairs of fine-tuned models and their base models, as well as for randomly selected model pairs. Summary statistics for these distributions are reported in Table 1(b). The figure shows that KL divergence after fine-tuning tends to be slightly smaller, as indicated by the left-skewed distribution. As noted in Oyama et al. (2025), models of the same type tend to have smaller KL divergence. To more accurately assess the effect of fine-tuning, we also analyze random model pairs sampled within the same model type.
Appendix F Details of Section 5
Figure 19: Re-rendering of the top panel of Fig. 6. Colors indicate mean log-likelihood, clipped at the bottom 5% across all values to improve visibility.
Figure 20: 3D t-SNE visualization of the left panel of Fig. 7. Deeper layers are represented with lighter colors, and the final layer is indicated by a white square.
Here, we describe the language models used in Section 5. For the trajectories shown in Fig. 6 and Fig. 19, we selected the 50 most downloaded models from those analyzed in Oyama et al. (2025). For specified model type analysis, we selected models with 32 layers from the llama-2, llama-3, and mistral types. For each of these types, we used the 10 most downloaded models. These models were used in the analysis shown in Fig. 7 and Fig. 20. A complete list of the models is provided in Table LABEL:tab:model_list_layer in Appendix G.
Fig. 19 shows the trajectories of models across layers, color-coded by mean log-likelihood, depending on layer depth. The mean log-likelihood increases as the layer depth increases. This indicates that the direction of trajectory progression corresponds to deeper layers, that is, to increasing mean log-likelihood. Additionally, Fig. 20 presents a 3D version of the left panel of Fig. 6. This illustration implies that the jump observed in Fig. 7 is not an artifact of the 2D t-SNE visualization, but rather results from a genuinely large KL divergence between adjacent layers.
Appendix G Model List
The models used in Sections 3, 4, and 5 are listed in Tables LABEL:tab:model_list_pretraining, LABEL:tab:model_list_ft, and LABEL:tab:model_list_layer, respectively. Table LABEL:tab:model_list_pretraining is sorted by model size; Table LABEL:tab:model_list_ft is sorted alphabetically by model name, considering the model generation, model type, and corresponding base model; and Table LABEL:tab:model_list_layer is sorted in descending order of download counts.
BibTeX entries for each model were selected according to the procedure described by Oyama et al. (2025). Note that, for the Pythia models trained with different random seeds van der Wal et al. (2025) and used in Section 3, the corresponding BibTeX entries in Table LABEL:tab:model_list_pretraining were prepared manually.
Table 5: List of 14 models in the experiments for pretraining models. “ID” denotes the index sorted by model size; “Model Name” denotes the name of the model.
ID Model Name
1 EleutherAI/pythia-410m Gao et al. (2020); Biderman et al. (2022, 2023)
2 EleutherAI/pythia-410m-seed1 van der Wal et al. (2025)
3 EleutherAI/pythia-410m-seed2 van der Wal et al. (2025)
4 EleutherAI/pythia-410m-seed3 van der Wal et al. (2025)
5 EleutherAI/pythia-410m-seed4 van der Wal et al. (2025)
6 EleutherAI/pythia-410m-seed5 van der Wal et al. (2025)
7 EleutherAI/pythia-410m-seed6 van der Wal et al. (2025)
8 EleutherAI/pythia-410m-seed7 van der Wal et al. (2025)
9 EleutherAI/pythia-410m-seed8 van der Wal et al. (2025)
10 EleutherAI/pythia-410m-seed9 van der Wal et al. (2025)
11 EleutherAI/pythia-1b Gao et al. (2020); Biderman et al. (2022, 2023)
12 EleutherAI/pythia-1.4b Gao et al. (2020); Biderman et al. (2022, 2023)
13 EleutherAI/pythia-2.8b Gao et al. (2020); Biderman et al. (2022, 2023)
14 EleutherAI/pythia-6.9b Gao et al. (2020); Biderman et al. (2022, 2023)
Table 6: List of 87 models in the experiments for fine-tuning models. “ID” denotes the alphabetical index considering the generation and the base model; “Model Name” denotes the name of the model; “Base Model Name” denotes the name of the base model before training; “Model Type” denotes the model classification defined in Oyama et al. (2025); “Generation” denotes the model generation in the training process; “DLs” denotes the total number of downloads.
ID Model Name Base Model Name Model Type Generation DLs
1 deepseek-ai/DeepSeek-Prover-V1.5-Base Xin et al. (2024) deepseek 1 230
2 tiiuae/falcon-rw-1b Brown et al. (2020); Press et al. (2022); Dao et al. (2022); Penedo et al. (2023) falcon 1 22921
3 google/gemma-1.1-7b-it Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Sakaguchi et al. (2019); Bisk et al. (2019); Chollet (2019); Zellers et al. (2019); Hendrycks et al. (2021); Chen et al. (2021); Austin et al. (2021); Cobbe et al. (2021); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 1 19835
4 google/gemma-7b Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Bisk et al. (2019); Chollet (2019); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Zellers et al. (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); Dhamala et al. (2021); Cobbe et al. (2021); Lin et al. (2022); Hartvigsen et al. (2022); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Dettmers et al. (2023); Gemini Team et al. (2024) gemma 1 72047
5 openchat/openchat-3.5-0106-gemma Wang et al. (2024a) gemma 1 7817
6 openlm-research/open_llama_3b Together Computer (2023); Touvron et al. (2023a); Geng and Liu (2023) llama-1 1 157405
7 codellama/CodeLlama-13b-Instruct-hf Rozière et al. (2024) llama-2 1 25248
8 meta-llama/Llama-2-13b-hf Touvron et al. (2023b) llama-2 1 120871
9 meta-llama/Llama-2-7b-hf Touvron et al. (2023b) llama-2 1 1294737
10 meta-llama/Meta-Llama-3-8B-Instruct AI@Meta (2024) llama-3 1 2058011
11 meta-llama/Meta-Llama-3-8B AI@Meta (2024) llama-3 1 675850
12 NousResearch/Hermes-2-Pro-Llama-3-8B "Teknium" et al. (2024b) llama-3 1 26797
13 mistralai/Mistral-7B-Instruct-v0.2 Jiang et al. (2023) mistral 1 3565248
14 mistralai/Mistral-7B-v0.1 Jiang et al. (2023) mistral 1 1750089
15 mlabonne/Marcoro14-7B-slerp mistral 1 3792
16 mlabonne/NeuralMonarch-7B mistral 1 13486
17 stabilityai/japanese-stablelm-base-gamma-7b Jiang et al. (2023) mistral 1 2072
18 beomi/KoRWKV-6B others 1 2127
19 microsoft/Orca-2-13b Mitra et al. (2023) others 1 13616
20 upstage/SOLAR-10.7B-v1.0 Kim et al. (2024b) others 1 24478
21 01-ai/Yi-1.5-9B 01. AI et al. (2025) others 1 20252
22 deepseek-ai/DeepSeek-Prover-V1.5-SFT Xin et al. (2024) deepseek-ai/DeepSeek-Prover-V1.5-Base deepseek 2 6459
23 euclaise/falcon_1b_stage1 tiiuae/falcon-rw-1b falcon 2 2119
24 lemon-mint/gemma-ko-7b-instruct-v0.71 google/gemma-1.1-7b-it gemma 2 2232
25 lemon-mint/gemma-7b-openhermes-v0.80 google/gemma-1.1-7b-it gemma 2 4867
26 google/gemma-7b-it Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Bisk et al. (2019); Chollet (2019); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Zellers et al. (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); Dhamala et al. (2021); Cobbe et al. (2021); Lin et al. (2022); Hartvigsen et al. (2022); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) google/gemma-7b gemma 2 66776
27 Telugu-LLM-Labs/Indic-gemma-7b-finetuned-sft-Navarasa-2.0 google/gemma-7b gemma 2 2025
28 lemon-mint/gemma-ko-7b-instruct-v0.62 openchat/openchat-3.5-0106-gemma gemma 2 7543
29 rinna/youri-7b Andonian et al. (2021); Touvron et al. (2023b); Zhao et al. (2023); Sawada et al. (2024) meta-llama/Llama-2-7b-hf llama-2 2 2512
30 aaditya/Llama3-OpenBioLLM-8B Singhal et al. (2022); Nori et al. (2023); Singhal et al. (2023); Pal and Sankarasubbu (2024b, a); Rafailov et al. (2024) meta-llama/Meta-Llama-3-8B llama-3 2 9308
31 cognitivecomputations/dolphin-2.9-llama3-8b meta-llama/Meta-Llama-3-8B llama-3 2 130977
32 cognitivecomputations/dolphin-2.9.1-llama-3-8b meta-llama/Meta-Llama-3-8B llama-3 2 7575
33 DeepMount00/Llama-3-8b-Ita meta-llama/Meta-Llama-3-8B llama-3 2 179401
34 dfurman/Llama-3-8B-Orpo-v0.1 meta-llama/Meta-Llama-3-8B llama-3 2 5146
35 johnsnowlabs/JSL-MedLlama-3-8B-v2.0 meta-llama/Meta-Llama-3-8B llama-3 2 11813
36 jondurbin/bagel-8b-v1.0 meta-llama/Meta-Llama-3-8B llama-3 2 7960
37 openchat/openchat-3.6-8b-20240522 Wang et al. (2024a) meta-llama/Meta-Llama-3-8B llama-3 2 10464
38 rinna/llama-3-youko-8b Andonian et al. (2021); Mitsuda et al. (2024); AI@Meta (2024); Sawada et al. (2024) meta-llama/Meta-Llama-3-8B llama-3 2 1468
39 ruslanmv/Medical-Llama3-8B meta-llama/Meta-Llama-3-8B llama-3 2 5457
40 lightblue/suzume-llama-3-8B-multilingual Devine (2024) meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 16442
41 MaziyarPanahi/Llama-3-8B-Instruct-v0.4 meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 1407
42 PathFinderKR/Waktaverse-Llama-3-KO-8B-Instruct AI@Waktaverse (2024); AI@Meta (2024) meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 2305
43 shenzhi-wang/Llama3-8B-Chinese-Chat Wang et al. (2024b) meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 55718
44 SJ-Donald/llama3-passthrough-chat meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 2241
45 swap-uniba/LLaMAntino-3-ANITA-8B-Inst-DPO-ITA Basile et al. (2023); AI@Meta (2024); Polignano et al. (2024) meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 5995
46 NousResearch/Hermes-2-Theta-Llama-3-8B "Teknium" et al. (2024a) NousResearch/Hermes-2-Pro-Llama-3-8B llama-3 2 12392
47 vicgalle/Configurable-Hermes-2-Pro-Llama-3-8B Gallego (2024) NousResearch/Hermes-2-Pro-Llama-3-8B llama-3 2 10273
48 abacusai/bigstral-12b-32k mistralai/Mistral-7B-Instruct-v0.2 mistral 2 5216
49 Mihaiii/Metis-0.3 mistralai/Mistral-7B-Instruct-v0.2 mistral 2 1279
50 alignment-handbook/zephyr-7b-sft-full mistralai/Mistral-7B-v0.1 mistral 2 11605
51 cognitivecomputations/dolphin-2.2.1-mistral-7b mistralai/Mistral-7B-v0.1 mistral 2 7292
52 cypienai/cymist-2-v02-SFT Lacoste et al. (2019) mistralai/Mistral-7B-v0.1 mistral 2 2717
53 HuggingFaceH4/mistral-7b-sft-beta mistralai/Mistral-7B-v0.1 mistral 2 7408
54 HuggingFaceH4/zephyr-7b-alpha Ding et al. (2023); Tunstall et al. (2023); Cui et al. (2024); Rafailov et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 12911
55 HuggingFaceH4/zephyr-7b-beta Ding et al. (2023); Tunstall et al. (2023); Cui et al. (2024); Rafailov et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 294019
56 ibm/merlinite-7b Sudalairaj et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 11156
57 INSAIT-Institute/BgGPT-7B-Instruct-v0.2 mistralai/Mistral-7B-v0.1 mistral 2 3265
58 Intel/neural-chat-7b-v3-1 Mukherjee et al. (2023) mistralai/Mistral-7B-v0.1 mistral 2 3976
59 kaist-ai/mistral-orpo-capybara-7k Hong et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 4821
60 Locutusque/Hercules-2.5-Mistral-7B mistralai/Mistral-7B-v0.1 mistral 2 1953
61 mistralai/Mistral-7B-Instruct-v0.1 Jiang et al. (2023) mistralai/Mistral-7B-v0.1 mistral 2 1370245
62 NousResearch/Hermes-2-Pro-Mistral-7B "interstellarninja" et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 14586
63 NousResearch/Nous-Hermes-2-Mistral-7B-DPO "Teknium" et al. (2024c) mistralai/Mistral-7B-v0.1 mistral 2 8725
64 openchat/openchat-3.5-0106 Wang et al. (2024a); OpenAI et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 26858
65 openchat/openchat-3.5-1210 Wang et al. (2024a); OpenAI et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 2123
66 statking/zephyr-7b-sft-full-orpo mistralai/Mistral-7B-v0.1 mistral 2 2278
67 teknium/OpenHermes-2.5-Mistral-7B mistralai/Mistral-7B-v0.1 mistral 2 100997
68 mlabonne/NeuralMarcoro14-7B mlabonne/Marcoro14-7B-slerp mistral 2 2015
69 mlabonne/AlphaMonarch-7B mlabonne/NeuralMonarch-7B mistral 2 12804
70 augmxnt/shisa-gamma-7b-v1 stabilityai/japanese-stablelm-base-gamma-7b mistral 2 151426
71 beomi/KoAlpaca-KoRWKV-6B beomi/KoRWKV-6B others 2 2288
72 Nexusflow/NexusRaven-V2-13B Nexusflow.ai team (2023); Rozière et al. (2024) codellama/CodeLlama-13b-Instruct-hf others 2 3966
73 haoranxu/ALMA-13B-Pretrain Xu et al. (2024a, b) meta-llama/Llama-2-13b-hf others 2 5510
74 ruslanmv/ai-medical-model-32bit meta-llama/Meta-Llama-3-8B-Instruct others 2 2810
75 Locutusque/Orca-2-13b-SFT-v4 microsoft/Orca-2-13b others 2 2345
76 Locutusque/Orca-2-13b-SFT-v6 microsoft/Orca-2-13b others 2 2319
77 Locutusque/Orca-2-13b-SFT_v5 microsoft/Orca-2-13b others 2 2191
78 mwitiderrick/open_llama_3b_code_instruct_0.1 openlm-research/open_llama_3b others 2 1252
79 NousResearch/Nous-Hermes-2-SOLAR-10.7B upstage/SOLAR-10.7B-v1.0 others 2 9918
80 upstage/SOLAR-10.7B-Instruct-v1.0 Kim et al. (2024b, a) upstage/SOLAR-10.7B-v1.0 others 2 67725
81 cognitivecomputations/dolphin-2.9.1-yi-1.5-9b 01-ai/Yi-1.5-9B others 2 4880
82 deepseek-ai/DeepSeek-Prover-V1.5-RL Xin et al. (2024) deepseek-ai/DeepSeek-Prover-V1.5-SFT deepseek 3 12223
83 euclaise/falcon_1b_stage2 euclaise/falcon_1b_stage1 falcon 3 4016
84 MaziyarPanahi/Llama-3-8B-Instruct-v0.8 MaziyarPanahi/Llama-3-8B-Instruct-v0.4 llama-3 3 7281
85 argilla/notus-7b-v1 alignment-handbook/zephyr-7b-sft-full mistral 3 7660
86 Intel/neural-chat-7b-v3-3 Yu et al. (2024) Intel/neural-chat-7b-v3-1 mistral 3 166226
87 MaziyarPanahi/Llama-3-8B-Instruct-v0.9 MaziyarPanahi/Llama-3-8B-Instruct-v0.8 llama-3 4 6241
Table 7: List of 59 models in the experiments for layer-wise models. “ID” denotes the index sorted in descending order of downloads; “Model Name” denotes the name of the model; “Model Type” denotes the model classification defined in Oyama et al. (2025); “Layers” denotes the number of layers in the model; “DLs” denotes the total number of downloads.
ID Model Name Model Type Layers DLs
1 mistralai/Mistral-7B-Instruct-v0.2 Jiang et al. (2023) mistral 32 3565248
2 meta-llama/Meta-Llama-3-8B-Instruct AI@Meta (2024) llama-3 32 2058011
3 mistralai/Mistral-7B-v0.1 Jiang et al. (2023) mistral 32 1750089
4 mistralai/Mistral-7B-v0.3 mistral 32 1443065
5 meta-llama/Llama-2-7b-chat-hf Touvron et al. (2023b) llama-2 32 1402244
6 mistralai/Mistral-7B-Instruct-v0.1 Jiang et al. (2023) mistral 32 1370245
7 meta-llama/Llama-2-7b-hf Touvron et al. (2023b) llama-2 32 1294737
8 TinyLlama/TinyLlama-1.1B-Chat-v1.0 others 22 1078500
9 TinyLlama/TinyLlama-1.1B-intermediate-step-1431k-3T others 22 798206
10 meta-llama/Meta-Llama-3-8B AI@Meta (2024) llama-3 32 675850
11 OpenAssistant/oasst-sft-4-pythia-12b-epoch-3.5 gpt_neox 36 458718
12 bigcode/starcoder2-3b Beltagy et al. (2020); Dao et al. (2022); Bavarian et al. (2022); Ainslie et al. (2023); Lozhkov et al. (2024) others 30 445316
13 lmsys/vicuna-7b-v1.5 Touvron et al. (2023b); Zheng et al. (2023) llama-2 32 368410
14 HuggingFaceH4/zephyr-7b-beta Ding et al. (2023); Tunstall et al. (2023); Cui et al. (2024); Rafailov et al. (2024) mistral 32 294019
15 meta-llama/Llama-2-13b-chat-hf Touvron et al. (2023b) llama-2 40 266090
16 google/gemma-2b Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Bisk et al. (2019); Chollet (2019); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Zellers et al. (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); Dhamala et al. (2021); Cobbe et al. (2021); Lin et al. (2022); Hartvigsen et al. (2022); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 18 256798
17 EleutherAI/gpt-neo-1.3B Gao et al. (2020); Black et al. (2021) others 24 243332
18 EleutherAI/gpt-j-6b Gao et al. (2020); Wang and Komatsuzaki (2021); Wang (2021); Su et al. (2023) gptj 28 241435
19 microsoft/phi-2 others 32 237268
20 EleutherAI/gpt-neo-2.7B Gao et al. (2020); Black et al. (2021) others 32 205503
21 huggyllama/llama-7b llama-1 32 192383
22 tiiuae/falcon-7b-instruct Shazeer (2019); Brown et al. (2020); Dao et al. (2022); Su et al. (2023); Almazrouei et al. (2023); Penedo et al. (2023) falcon 32 179952
23 DeepMount00/Llama-3-8b-Ita llama-3 32 179401
24 Intel/neural-chat-7b-v3-3 Yu et al. (2024) mistral 32 166226
25 openlm-research/open_llama_3b Together Computer (2023); Touvron et al. (2023a); Geng and Liu (2023) llama-1 26 157405
26 augmxnt/shisa-gamma-7b-v1 mistral 32 151426
27 facebook/opt-1.3b Brown et al. (2020); Zhang et al. (2022) opt 24 139758
28 Qwen/Qwen1.5-1.8B Bai et al. (2023) others 24 137256
29 codellama/CodeLlama-7b-Instruct-hf Rozière et al. (2024) llama-2 32 133523
30 cognitivecomputations/dolphin-2.9-llama3-8b llama-3 32 130977
31 Qwen/Qwen1.5-7B Bai et al. (2023) others 32 122768
32 meta-llama/Llama-2-13b-hf Touvron et al. (2023b) llama-2 40 120871
33 microsoft/phi-1_5 Li et al. (2023) others 24 112476
34 NousResearch/Meta-Llama-3-8B-Instruct AI@Meta (2024) llama-3 32 104388
35 tiiuae/falcon-7b Shazeer (2019); Brown et al. (2020); Gao et al. (2020); Dao et al. (2022); Su et al. (2023); Almazrouei et al. (2023); Penedo et al. (2023) falcon 32 104010
36 google/gemma-2b-it Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Bisk et al. (2019); Chollet (2019); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Zellers et al. (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); Dhamala et al. (2021); Cobbe et al. (2021); Lin et al. (2022); Hartvigsen et al. (2022); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 18 103851
37 teknium/OpenHermes-2.5-Mistral-7B mistral 32 100997
38 mosaicml/mpt-7b-chat Henry et al. (2020); Press et al. (2022); Dao et al. (2022); MosaicML NLP Team (2023) others 32 88069
39 Qwen/CodeQwen1.5-7B-Chat Bai et al. (2023) others 32 77169
40 GritLM/GritLM-7B Muennighoff et al. (2025) mistral 32 72991
41 google/gemma-7b Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Bisk et al. (2019); Chollet (2019); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Zellers et al. (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); Dhamala et al. (2021); Cobbe et al. (2021); Lin et al. (2022); Hartvigsen et al. (2022); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Dettmers et al. (2023); Gemini Team et al. (2024) gemma 28 72047
42 upstage/SOLAR-10.7B-Instruct-v1.0 Kim et al. (2024b, a) others 48 67725
43 google/gemma-7b-it Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Bisk et al. (2019); Chollet (2019); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Zellers et al. (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); Dhamala et al. (2021); Cobbe et al. (2021); Lin et al. (2022); Hartvigsen et al. (2022); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 28 66776
44 codellama/CodeLlama-7b-hf Rozière et al. (2024) llama-2 32 66040
45 facebook/opt-2.7b Brown et al. (2020); Zhang et al. (2022) opt 32 61450
46 shenzhi-wang/Llama3-8B-Chinese-Chat Wang et al. (2024b) llama-3 32 55718
47 VAGOsolutions/Llama-3-SauerkrautLM-8b-Instruct llama-3 32 55400
48 lmsys/vicuna-13b-v1.5 Touvron et al. (2023b); Zheng et al. (2023) llama-2 40 48653
49 Qwen/Qwen2-1.5B Yang et al. (2024) others 28 46510
50 openlm-research/open_llama_7b Together Computer (2023); Touvron et al. (2023a); Geng and Liu (2023) llama-1 32 44968
51 LinkSoul/Chinese-Llama-2-7b llama-2 32 40799
52 mistral-community/Mistral-7B-v0.2 mistral 32 30074
53 NousResearch/Hermes-2-Pro-Llama-3-8B "Teknium" et al. (2024b) llama-3 32 26797
54 lightblue/suzume-llama-3-8B-multilingual Devine (2024) llama-3 32 16442
55 abacusai/Llama-3-Smaug-8B Pal et al. (2024) llama-3 32 12873
56 NousResearch/Nous-Hermes-llama-2-7b llama-2 32 12019
57 togethercomputer/LLaMA-2-7B-32K llama-2 32 8884
58 elyza/ELYZA-japanese-Llama-2-7b-instruct Touvron et al. (2023b); Sasaki et al. (2023) llama-2 32 6296
59 togethercomputer/Llama-2-7B-32K-Instruct Liu et al. (2023) llama-2 32 5596

### v2

@@
-Revealing Language Model Trajectories via Kullback-Leibler Divergence
+Establishing a Scale for Kullback–Leibler Divergence in Language Models Across Various Settings
 Abstract
-A recently proposed method enables efficient estimation of the KL divergence between language models, including models with different architectures, by assigning coordinates based on log-likelihood vectors.
-To better understand the behavior of this metric, we systematically evaluate KL divergence across a wide range of conditions using publicly available language models.
-Our analysis covers comparisons between pretraining checkpoints, fine-tuned and base models, and layers via the logit lens.
-We find that trajectories of language models, as measured by KL divergence, exhibit a spiral structure during pretraining and thread-like progressions across layers.
-Furthermore, we show that, in terms of diffusion exponents, model trajectories in the log-likelihood space are more constrained than those in weight space.
-Revealing Language Model Trajectories via Kullback-Leibler Divergence
-Ryo Kishino1  Yusuke Takase1  Momose Oyama1,2 Hiroaki Yamagiwa1   Hidetoshi Shimodaira1,2 1Kyoto University 2RIKEN kishino.ryo.32s@st.kyoto-u.ac.jp, {y.takase, oyama.momose}@sys.i.kyoto-u.ac.jp, {h.yamagiwa, shimo}@i.kyoto-u.ac.jp
+Log-likelihood vectors define a common space for comparing language models as probability distributions, enabling unified comparisons across heterogeneous settings.
+We extend this framework to training checkpoints and intermediate layers, and establish a consistent scale for KL divergence across pretraining, model size, random seeds, quantization, fine-tuning, and layers.
+Analysis of Pythia pretraining trajectories further shows that changes in log-likelihood space are much smaller than in weight space, resulting in subdiffusive learning trajectories and early stabilization of language-model behavior despite weight drift.
+Establishing a Scale for Kullback–Leibler Divergence
+in Language Models Across Various Settings
+Ryo Kishino1   Yusuke Takase1   Momose Oyama1,2 Hiroaki Yamagiwa1   Hidetoshi Shimodaira1,2 1Kyoto University 2RIKEN kishino.ryo.32s@st.kyoto-u.ac.jp, {oyama.momose, y.takase}@sys.i.kyoto-u.ac.jp, {h.yamagiwa, shimo}@i.kyoto-u.ac.jp
 1 Introduction
-Figure 1: Visualization of the training trajectories of Pythia models using t-SNE applied to log-likelihood vectors.
+To understand learning dynamics and intermediate-layer representations in language models, it is essential to quantify changes in model behavior and compare them across models.
+While traditional analyses rely on weight parameters Chen et al.
+(2022); Kunin et al.
+(2024); Singh et al.
+(2025); Gigante et al.
+(2019), weight permutation symmetries Hecht-Nielsen (1990) and architectural dependencies hinder direct comparisons between models with different learning methods or designs.
+Figure 1: Visualization of the pretraining trajectories of Pythia models across five model sizes and seven random seeds using t-SNE (perplexity=30) applied to log-likelihood vectors.
 Each point represents a model checkpoint, with lighter colors indicating earlier steps.
-The thickness of the connecting lines reflects the magnitude of KL divergence between checkpoints.
-The top panel shows trajectories for five model sizes and multiple random seeds.
-The bottom panel shows only the five sizes, with both 2D and 3D t-SNE visualizations for clarity.
-See Section 3 for details.
+The thickness of the connecting lines reflects the magnitude of KL divergence between successive checkpoints.
+See Appendix D for additional visual analyses.
+Oyama et al.
+(2025) proposed the log-likelihood vector, a common representation of a model’s probabilities over a set of texts.
+They showed that models with different architectures can be embedded in the same space.
+Moreover, since the squared Euclidean distance in this space estimates the Kullback–Leibler (KL) divergence, model comparison can be formulated as a geometric problem.
+In this study, we demonstrate that log-likelihood vectors provide a consistent representation not only for fully trained models but also for training checkpoints, quantized models, and intermediate layers.
+We therefore handle diverse models within a single coordinate system (see Figs.
+1 and 2).
+We establish KL divergence scales across various settings, including checkpoints, model sizes, random seeds, quantization, fine-tuning, and layers as practical metrics for model comparison.
+Figure 2: (a) t-SNE visualization (perplexity=30) of the original models and their 8-bit quantized models.
+(b) t-SNE visualization (perplexity=20) of the base models and their fine-tuned models.
+Lines represent parent-child relationships.
+Generation refers to the depth from the root model within each lineage.
+(c) t-SNE visualization (perplexity=30) of the trajectories across layers for 50 models, obtained by applying the logit lens to each layer.
+Deeper layers are represented with lighter colors, and the final layer is indicated by a white square.
+Moreover, experiments on Pythia show that when focusing on pretraining trajectories and estimating KL-scaling exponents, large changes in weight space remain remarkably small in log-likelihood space.
+This discrepancy appears as subdiffusive behavior in later training steps, indicating that the model’s output is far more stable.
+2 Preliminaries: Model Map and KL
+Log-likelihood vector.
+The probability that a language model $p$ generates a text (i.e., a sequence of tokens) $x=(y_{0},\ldots,y_{M})$ is given by $p(x)=\prod_{m=1}^{M}p(y_{m}\mid y_{0},\ldots,y_{m-1})$.
+Oyama et al.
+(2025) defined the log-likelihood vector of a model $p$ over a predefined text set [^1] Olsson et al.
+(2022) used token-level log-likelihoods.
+$\{x_{1},\ldots,x_{N}\}$ as
+$\bm{\ell}=(\log p(x_{1}),\ldots,\log p(x_{N}))^{\top}\in\mathbb{R}^{N},$
+and showed that it is a useful feature representation for model comparison and downstream performance prediction. This method has also been used in subsequent work such as Harada et al. (2025).
+KL divergence.
+Let $\{p_{i}\}_{i=1}^{K}$ be a set of language models, and $\bm{L}=(\bm{\ell}_{1},\ldots,\bm{\ell}_{K})^{\top}\in\mathbb{R}^{K\times N}$ be the matrix formed by stacking the log-likelihood vectors $\bm{\ell}_{i}$ of each model.
+Define $\bm{Q}=(\bm{q}_{1},\ldots,\bm{q}_{K})^{\top}$ as the matrix obtained by double-centering [^2] Row-wise centering (over texts) and column-wise centering (over models), respectively, of the log-likelihood matrix.
+$\bm{L}$.
+Then, the KL divergence between two models $p_{i}$ and $p_{j}$ can be approximated as
+$2\mathrm{KL}(p_{i},p_{j})\approx\|\bm{q}_{i}-\bm{q}_{j}\|^{2}/N.$
+In other words, by associating each model $p_{i}$ with a coordinate $\bm{q}_{i}\in\mathbb{R}^{N}$, KL divergence is approximated by squared Euclidean distance in this space.
+Oyama et al.
+(2025) referred to this coordinate system as the model map.
+For the interpretation of the KL divergence, the standard error, and the dependence on the text set, see Appendix C.
+Model map for various settings.
+Log-likelihood vectors can be computed not only for fully trained language models but also for models during training (Figs.
+1 and 2b) and for quantized models (Fig.
+2a).
+Moreover, by treating the network up to each layer as a single model via the logit lens nostalgebraist (2020), intermediate layers can also be analyzed (Fig.
+2c).
+These analyses are difficult to carry out with weight-based approaches Gigante et al.
+(2019); Chen et al.
+(2022); Kunin et al.
+(2024); Singh et al.
+(2025).
+General experimental settings.
+For the text set, we use the same 10,000 texts extracted from the Pile arxiv:2101.00027 as in Oyama et al.
+(2025).
+The values of KL divergence are normalized by the average text length, $\bar{B}=972.3188$ bytes, and are reported in units of bits per byte.
+Accordingly, the matrix $\bm{Q}$ is divided in advance by $\sqrt{2N\bar{B}\log 2}$, so that the squared Euclidean distance between rows of $\bm{Q}$ directly approximates the KL divergence in bits per byte.
+Figure 3: Scale of KL divergence (bits per byte) across various settings.
+The first four entries from the left correspond to pretraining-related comparisons, showing the KL divergence between consecutive saved checkpoints at 1k-step intervals during the last 50k steps of training, at 1k-step intervals during the first 50k steps, across different random seeds at the final checkpoint, and across different model sizes.
+The subsequent entries show the KL divergence before and after quantization (8-bit and 4-bit), before and after fine-tuning, between randomly selected pairs within the same model type, between randomly selected pairs across all model types, and between adjacent layers.
+The vertical axis is shown on a symlog scale.
+See Appendix A and Table 1 for more details.
+3 KL Divergence Across Various Settings
+In this section, we measure and summarize KL divergence under a variety of settings, and present the results in Fig.
+3.
+KL divergence is an interpretable quantity on an absolute scale; for example, a value of 0.1 bits/byte corresponds to only 0.1 bit per 8-bit text unit, whereas 10 bits/byte indicates a very large difference, exceeding the 8-bit scale of the text itself [^3] Since KL divergence is unbounded, it can exceed 8 bits/byte despite the 8-bit representation of text..
+As a reference point, the effective per-byte entropy of the text set in our experiments is approximately 0.71 bits/byte (see Appendix C).
+Our experiments reveal that each setting is associated with a characteristic scale of KL divergence.
+Even the same value of 0.1 bits/byte can be interpreted as a substantial change between consecutive saved 1k-step checkpoints during training, but as a minor difference when comparing models of different types.
+3.1 Pretraining
+Settings.
+We use publicly available pretraining checkpoints from the Pythia suite arxiv:2304.01373 for model sizes 410M, 1B, 1.4B, 2.8B, and 6.9B.
+For the 410M model, checkpoints are available for nine random seeds (1–9) released by wal2025polypythias.
+From these, we exclude seeds 3 and 4 due to training loss spikes detected via KL divergence, leaving seven seeds for our experiments.
+Checkpoints are saved at training steps 0, 1, 2, 4, …, 512, 1k, 2k, 3k, …, and 143k.
+We exclude the 1B checkpoint at step 116k due to anomalous weights and remove texts exhibiting irregular log-likelihood variations across checkpoints.
+Details of the preprocessing, data filtering, and anomaly detection procedures for pretraining are provided in Appendix E.
+KL divergence.
+For all model sizes, the KL divergence between consecutive saved checkpoints at 1k-step intervals after the warmup phase is typically 0.05–0.1 bits/byte during the early stage of training, and becomes extremely small in the later stage, ranging from about 0.01 to 0.05 bits/byte.
+At the final checkpoint, the KL divergence across different random seeds for the 410M model is around 0.1 bits/byte, while that across different model sizes ranges from 0.15 to 1.7 bits/byte, indicating substantially larger differences.
+3.2 Quantization
+Settings.
+We analyze the impact of post-training quantization Dettmers et al.
+(2022).
+We use a subset of the 1,018 language models analyzed in Oyama et al.
+(2025), selecting the 50 most downloaded models (Fig.
+2a); see Appendix F for details on the models used.
+KL divergence.
+The median KL divergence before and after 8-bit quantization is 0.44 bits/byte, while it is slightly larger for 4-bit quantization at 0.49 bits/byte.
+As shown in Fig.
+3, the distributions exhibit low variance in both cases, indicating that the degree of degradation induced by quantization is relatively consistent across models.
+Furthermore, in Fig.
+2a, the shifts induced by quantization appear to be aligned in both direction and magnitude within each model type, suggesting that quantization acts as a structured perturbation in the log-likelihood space rather than as random noise.
+3.3 Fine-tuning Effects
+Settings.
+Out of the 1,018 language models analyzed in Oyama et al.
+(2025), we use 66 pairs of fine-tuned and base models (87 models in total; Fig.
+2b).
+The parent-child relationships are identified using Hugging Face Hub API.
+See Appendix G for details.
+KL divergence.
+Fig.
+3 shows that fine-tuning induces a relatively small but non-negligible change within a model type, smaller than the variation among randomly paired models of the same type and much smaller than the variation across model types; the corresponding median KL divergence values are 0.40, 0.95, and 2.2 bits/byte.
+3.4 Across Layers
+Settings.
+We investigate model trajectories across layers by treating the network up to each layer as an individual model using the logit lens nostalgebraist (2020).
+We use the same set of 50 models as those used in Section 3.2 (Fig.
+2c).
+KL divergence.
+The KL divergence between adjacent layers spans a wide range, with a median of 3.0 bits/byte and a standard deviation of 13 bits/byte.
+Moreover, as shown in Fig.
+2c, models within the same type exhibit similar trajectory trends across layers.
+4 Scaling Analysis of KL Divergence Along Pretraining Trajectories
+In this section, we examine the scaling behavior of changes in KL divergence over training time along pretraining trajectories of Pythia models, using the same checkpoints as in Section 3.1.
+4.1 Trajectory on the Log-Likelihood Space
+Figure 4: KL divergence between consecutive saved checkpoints of Pythia during pretraining.
+(Left) Warmup phase with non-uniform checkpoint spacing.
+(Right) Post-warmup phase with checkpoints saved every 1k training steps.
+Fig.
+1 visualizes pretraining trajectories of Pythia models across five model sizes and seven random seeds in the log-likelihood space using t-SNE, enabling joint comparisons that are difficult in weight space [^4] Two-dimensional t-SNE visualizations may exhibit apparent jumps due to dimensionality constraints, whereas three-dimensional t-SNE visualizations confirm continuity in the original high-dimensional space; see Appendix D..
+Fig.
+4 shows the KL divergence between consecutive saved checkpoints during pretraining.
+After the warmup phase, the KL divergence between checkpoints saved every 1k steps decreases and, after step 10k, remains within a narrow range regardless of model size, whereas substantially larger values are observed during warmup.
+4.2 Comparison with Weight Space
+We compare the stability of training trajectories by estimating diffusion exponents that characterize the power-law scaling of KL divergence.
+According to Kunin et al.
+(2024), after model performance has converged, the weights diffuse following a power law with exponent $c_{\mathrm{w}}$.
+Specifically, with respect to an initial step $t_{0}$, the squared Euclidean distance satisfies $\|\bm{W}_{t}-\bm{W}_{t_{0}}\|^{2}\propto|t-t_{0}|^{c_{\mathrm{w}}}$.
+We show that an analogous power-law relationship also holds for the log-likelihood vectors $\bm{q}_{t}$, namely $\|\bm{q}_{t}-\bm{q}_{t_{0}}\|^{2}\propto|t-t_{0}|^{c_{\mathrm{q}}}$, which corresponds to the KL divergence.
+In Fig.
+5, the weights $\bm{W}$ exhibit Brownian-like diffusion with $c_{\mathrm{w}}\approx 1$, whereas the diffusion of $\bm{q}$ is substantially suppressed with $c_{\mathrm{q}}\approx 0.2$.
+Under a fractional Brownian motion interpretation, these exponents correspond to effective fractal dimensions of $D_{\mathrm{w}}\approx 2$ and $D_{\mathrm{q}}\approx 10$, respectively [^5] For diffusion exponent $c$, Brownian motion corresponds to $c=1$, while $c\neq 1$ indicates anomalous diffusion.
+If the process is fractional Brownian motion Mandelbrot and Van Ness (1968), the Hurst exponent is $H=c/2$, and the Hausdorff dimension of the trajectory is $D=1/H$ (Adler, 1981, Theorem 8.4.1)..
+In addition, the right panel of Fig.
+5 illustrates how the exponents $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ vary with the starting step $t_{0}$.
+Compared to weight trajectories, the scaling of KL divergence in log-likelihood space exhibits greater geometric complexity due to many-to-one folding, thus remaining confined to a narrow region and suggesting early stabilization of language-model behavior despite continued drift in weight space.
+The difference in dynamics between these two spaces may be attributed to folding induced by permutation symmetries in multilayer neural networks Fukumizu and Amari (2000) and to the flat minima hypothesis Keskar et al.
+(2017).
+A detailed discussion is provided in Appendix B.
+Figure 5: (Left) Temporal evolution of the squared Euclidean distance in the weights and log-likelihood vectors from steps 120k to 130k for Pythia.
+(Right) Diffusion exponent as a function of the starting point $t_{0}$.
+The exponent is estimated using the least squares method.
+5 Conclusion
+We measured KL divergence between language models across a wide range of conditions and established a unified and interpretable scale.
+We further examined the scaling behavior of changes in KL divergence along learning trajectories.
+Limitations
+  • We conduct our experiments using the same set of 10,000 text chunks as those used in Oyama et al.
+(2025), and the measured KL divergence therefore depends on the choice of the text set.
+However, we verified that different random subsets of 10,000 text chunks from the same Pile corpus yield highly consistent KL estimates (Appendix C), while the impact of domain shifts across text sets has not been examined.
+  • The set of models analyzed in this study is limited and does not fully cover the diversity of existing language models.
+In particular, pretraining trajectory analyses are conducted only on the Pythia family, and experiments involving fine-tuning, quantization, and layer-wise analysis rely on selected subsets of models due to data availability and metadata constraints, such as incomplete base-model information in model cards.
+  • In the layer-wise analysis, the logits obtained from shallow layers via the logit lens contain substantial noise.
+This issue could be addressed by using the tuned lens Belrose et al.
+(2023), an improved version of the logit lens.
+However, the number of publicly available pretrained tuned lenses is currently limited.
+We believe this limitation can be overcome in future work by training tuned lenses.
+  • The checkpoints for Pythia are available only at 1k-step intervals except for the early stages of training.
+Accordingly, our analysis focuses on trajectory dynamics at scales of 1k steps or larger, while finer-scale behavior below this resolution is not examined.
+This limitation does not affect our analysis of dynamics at coarser temporal scales.
+  • While we observe structured behaviors such as subdiffusion in the log-likelihood space, this work represents an initial attempt at trajectory analysis in this space, and our analysis of their underlying causes and implications for learning dynamics remains preliminary and has not yet been quantitatively established.
+Acknowledgments
+We thank Luis Iván Hernández Ruíz for helpful discussions.
+This work was partially supported by JSPS KAKENHI JP22H05106 and JP23H03355 (to HS), JST CREST JPMJCR21N3 (to HS), JSPS KAKENHI JP25K24366 (to HY), and JST BOOST JPMJBS2407 (to YT and MO).
+Appendix A Details of Fig. 3
+In Fig.
+3, the vertical axis, i.e., the KL divergence values, is plotted using a symmetric logarithmic scale (symlog).
+The scale is linear up to 0.1 bits/byte and logarithmically scaled beyond that point.
+For the color bar, values are clipped at 10 bits/byte.
+Detailed statistics corresponding to Fig.
+3 are presented in Table 1(a).
+In addition, for Pythia-410M, we report the KL divergence between checkpoints separated by 1k steps during pretraining, as well as the KL divergence across different seeds and model sizes at the final checkpoint, in Table 1(b).
+Furthermore, for the three concrete models, the KL divergence between selected layers is shown in Table 1(c).
+Median Mean SD
+Later 0.011 0.014 0.0050
+Early 0.067 0.078 0.047
+Seed 0.12 0.12 0.0071
+Size 0.48 0.61 0.47
+8-bit 0.44 0.80 1.9
+4-bit 0.49 0.91 2.0
+FT 0.40 1.5 3.6
+Intra-type 0.95 2.8 6.4
+Random 2.2 5.3 8.8
+Layer 3.0 8.1 13
+(a) Summary statistics (median, mean, SD) of KL divergence across various settings. This table corresponds to Fig. 3.
 between checkpoints between seeds between model sizes
-10$\to$11k 50$\to$51k 100$\to$101k avg. 1B 1.4B 2.8B 6.9B
+[…]
 410M 0.069 0.053 0.023 0.11 0.28 0.52 1.0 1.7
 ($\pm$0.0011) ($\pm$0.0010) ($\pm$0.00042) ($\pm$ 0.0024) ($\pm$0.010) ($\pm$0.018) ($\pm$0.036) ($\pm$0.060)
-(a) KL divergence for Pythia 410M, measured between checkpoints saved during pretraining, different random seeds, and different model sizes.
-For the inter-size comparison, the final checkpoints were used.
-See Section 3 for details.
-Type (#pair) Pair Median Mean SD
-mistral (25) fine-tuning 0.32 0.37 0.32
-random 0.79 0.74 0.39
-llama-3 (20) fine-tuning 0.20 0.67 0.80
-random 0.66 0.87 0.66
-all (66) fine-tuning 0.40 1.5 3.6
-random 2.1 5.0 9.5
-(b) Summary statistics (median, mean, SD) of KL divergence between fine-tuned and base models, and between random model pairs. See Section 4 for details.
-Model 1$\to$2 16$\to$17 31$\to$32
+(b) KL divergence for Pythia 410M, measured between checkpoints saved during pretraining, across different random seeds, and across different model sizes.
+For the inter-seed and inter-size comparisons, the final checkpoints are used.
+[…]
 Llama-2-7b-hf 0.61 1.6 1.3
@@
 (± 0.10) (± 0.12) (± 0.54)
-(c) KL divergence between adjacent layers within models, measured using the logit lens. See Section 5 for details.
+(c) KL divergence between adjacent layers within specific models, measured using the logit lens.
 Table 1: KL divergence in bits per byte between language models under various conditions.
@@
 See Appendix C for the formula used to compute the standard errors.
-The growing diversity of language models has created a need to quantify their differences systematically.
-Treating them as probability distributions over text, the Kullback-Leibler (KL) divergence serves as a natural comparison metric.
-Oyama et al.
-(2025) proposed using log-likelihood vectors on a fixed text set to assign model coordinates in a Euclidean space.
-In this space, the squared distance between models approximates their KL divergence, making model comparison a geometric problem.
-However, their analysis did not cover the typical KL divergence values.
-As an early report on model analysis leveraging log-likelihood vectors, we use publicly available models to measure the KL divergence between checkpoints saved during pretraining (Section 3), fine-tuned and base models (Section 4), and different layers (Section 5) as shown in Table 1.
-Furthermore, we analyze the trajectories of language models throughout training and across layers, yielding several notable observations.
-As shown in Fig.
-1, portions of the pretraining trajectories exhibit a spiral structure.
-Moreover, from the perspective of anomalous diffusion, model dynamics in the log-likelihood space become increasingly constrained during the later steps of training, compared to those in weight space.
-2 Review: Model Map
-Log-likelihood vector.
-The probability that a language model $p$ generates a text (i.e., a sequence of tokens) $x=(y_{0},\ldots,y_{M})$ is given by $p(x)=\prod_{m=1}^{M}p(y_{m}\mid y_{0},\ldots,y_{m-1})$.
-Oyama et al.
-(2025) defined the log-likelihood vector of a language model $p$ over a predefined text set $\{x_{1},\ldots,x_{N}\}$ as
-$\bm{\ell}=(\log p(x_{1}),\ldots,\log p(x_{N}))^{\top}\in\mathbb{R}^{N},$
-and showed that it serves as a useful feature representation for comparing models and predicting downstream task performance.
-KL divergence.
-Let $\{p_{i}\}_{i=1}^{K}$ be a set of language models, and $\bm{L}=(\bm{\ell}_{1},\ldots,\bm{\ell}_{K})^{\top}\in\mathbb{R}^{K\times N}$ be the matrix formed by stacking the log-likelihood vectors $\bm{\ell}_{i}$ of each model.
-Define $\bm{Q}=(\bm{q}_{1},\ldots,\bm{q}_{K})^{\top}$ as the matrix obtained by double-centering [^1] Row-wise centering followed by column-wise centering.
-$\bm{L}$.
-Then, the KL divergence between two models $p_{i}$ and $p_{j}$ can be approximated as
-$2\mathrm{KL}(p_{i},p_{j})\approx\|\bm{q}_{i}-\bm{q}_{j}\|^{2}/N.$
-In other words, by associating each model $p_{i}$ with a coordinate $\bm{q}_{i}\in\mathbb{R}^{N}$, KL divergence is approximated by squared Euclidean distance in this space.
-This coordinate system is referred to as the model map in Oyama et al.
+Appendix B Related Work and Discussions on Diffusion Exponents
+Anomalous diffusion in weight space.
+It has been reported that, during the training of neural networks, trajectories in weight space can exhibit anomalous diffusion, and this phenomenon has recently attracted considerable research interest Chen et al.
+(2022); Kunin et al.
+(2024); Ly and Gong (2025).
+These studies analyze training dynamics through the temporal scaling of distances in weight space and report deviations from standard Brownian motion.
+Diffusion exponents in weight and log-likelihood spaces.
+In Section 4, we estimated the diffusion exponents $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ for trajectories in weight space and log-likelihood space, respectively.
+Our results show that during the early phase of training, both exponents take relatively large values ($c_{\mathrm{w}}=1.2$ and $c_{\mathrm{q}}=0.8$), but decrease as training progresses.
+In the later stages, the exponent in weight space approaches $c_{\mathrm{w}}\approx 1$, corresponding to Brownian-like diffusion, whereas the exponent in log-likelihood space becomes much smaller ($c_{\mathrm{q}}\approx 0.2\ll 1$), which is characteristic of subdiffusion.
+In our discussion, fractional Brownian motion is used purely as an interpretive model to relate the observed diffusion exponents to effective fractal dimensions, rather than as a generative model of the training dynamics.
+We note that the estimated diffusion exponent $c_{\mathrm{q}}$ exhibits systematic variation across model sizes, as shown in the right panel of Fig.
+5.
+In particular, smaller models tend to show lower values ($c_{\mathrm{q}}\approx 0.1\text{--}0.2$), while larger models exhibit higher values ($c_{\mathrm{q}}\approx 0.2\text{--}0.3$).
+While we report $c_{\mathrm{q}}\approx 0.2$ as a representative value capturing the typical subdiffusive behavior in log-likelihood space, a more detailed analysis of model-size dependence is beyond the scope of this paper.
+Interpretation via loss landscape and flat minima.
+In our experimental setting, texts are sampled from a corpus that closely approximates the training data.
+Consequently, each component of the log-likelihood vector corresponds to a term contributing to the cross-entropy loss, and the negative average of these components serves as a reasonable empirical estimate of the expected loss.
+The marked contrast between the near-Brownian diffusion in weight space and the strongly suppressed diffusion in log-likelihood space indicates that, even as model weights continue to change, the output probability distributions remain relatively stable.
+This observation is consistent with the hypothesis that stochastic gradient descent tends to favor flat regions of the loss landscape Keskar et al.
 (2017).
-In particular, we refer to a sequence of models represented as coordinates in the model map as a trajectory.
-General experimental settings.
-For the text set, we use the same 10,000 texts extracted from the Pile Gao et al.
-(2020) as in Oyama et al.
-(2025) [^2] https://github.com/shimo-lab/modelmap/.
-The values of KL divergence are normalized by the average text length, $\bar{B}=972.3188$ bytes, and are reported in units of bits per byte.
-In particular, the matrix $\bm{Q}$ is divided by $\sqrt{2N\bar{B}\log 2}$ in advance, so that the squared Euclidean distance directly approximates the KL divergence in bits per byte.
-All visualizations are based on this $\bm{Q}$.
-3 Pretraining
-In the Pythia suite Biderman et al.
-(2023), the checkpoints saved during pretraining are publicly available for each model size.
-Additionally, van der Wal et al.
-(2025) released checkpoints for smaller Pythia models trained with different random seeds.
-In this section, we investigate how a language model’s trajectory evolves during pretraining by analyzing these checkpoints.
-Settings.
-We use checkpoints from the standard Pythia models with sizes 410M, 1B, 1.4B, 2.8B, and 6.9B.
-In addition, for the 410M model, we supplement our analysis with checkpoints from seven different random seeds.
-The checkpoints are saved at training steps 0, 1, 2, 4, …, 512, 1k, 2k, 3k, …, and 143k.
-We excluded the 1B checkpoint at step 116k due to anomalous weights, and removed texts with irregular log-likelihood variations across checkpoints.
-See Appendix D.1 for details.
-KL divergence.
-Table 1(a) shows the KL divergence for Pythia 410M.
-The KL divergence between checkpoints saved every 1k steps gradually decreases as training progresses, and after step 10k, it stays within a small range (approximately 0.02 to 0.07 bit/byte), as also shown in the right panel of Fig.
-4.
-Moreover, the values between different random seeds are around 0.1 bit/byte, while those between different sizes range from 0.3 to 2 bit/byte, indicating larger differences.
-See Appendix D.2 for details.
-Trajectory.
-Figure 2: Trajectory of Pythia 2.8B during pretraining, visualized by applying 3D t-SNE, followed by PCA on the later half portion. (Left) PC2 vs. PC3. (Right) Autocorrelation function of PC3.
-Model Size 410M 1B 1.4B 2.8B 6.9B
-Period (step) 30k 13k 15k 14k 12k
-Table 2: Period of the spiral trajectory of Pythia during pretraining. See Appendix D.3 for details.
-Fig.
-1 shows a simultaneous visualization of the pretraining trajectories of five model sizes using t-SNE van der Maaten and Hinton (2008) [^3] We used the scikit-learn implementation with the default perplexity value of 30., which is not possible with weight-based representations.
-The trajectories evolve continuously during training [^4] The apparent jumps in the 2D trajectories of the 2.8B and 6.9B models are artifacts of t-SNE, which tends to separate overlapping points.
-These trajectories appear continuous in 3D t-SNE visualizations..
-Moreover, the 3D visualization suggests that the trajectories exhibit a spiral structure.
-As shown in Fig.
-2, applying PCA to the later portion of the trajectory obtained via t-SNE reveals a circular pattern.
-The autocorrelation function of PC3 exhibits periodicity.
-The estimated periods of the spiral trajectories are shown in Table 2.
-Similar phenomena in weight trajectories have been studied previously Cohen et al.
-(2021); Netay (2024) described in detail in Appendix A.
-In addition, refer to Appendix D.4 for a trajectory visualization using PCA.
-Comparison with weights.
-Figure 3: Temporal evolution of the squared Euclidean distance in the weights and log-likelihood vectors from step 120k to step 130k for Pythia.
-The diffusion exponent $c$ is estimated using the least squares method.
-Figure 4: (Left) Diffusion exponent $c$ as a function of the starting point $t_{0}$.
-(Right) KL divergence between consecutive checkpoints of Pythia (corresponding to Table 1(a)).
-See Appendix D.5 for details.
-Differences between models with the same architecture can be measured by the distance between their weights.
-Let $\bm{W}_{t}$ be the weights of a neural network at step $t$.
-According to Kunin et al.
-(2024), after model performance has converged, the weights diffuse following a power law with exponent $c$ [^5] Brownian motion corresponds to $c=1$, while anomalous diffusion is defined as $c\neq 1$.
-Specifically, $c>1$ means superdiffusion, and $c<1$ means subdiffusion.
-If the evolution is fractional Brownian motion Mandelbrot and Van Ness (1968), the Hurst exponent corresponds to $H=c/2$..
-That is, with respect to an initial step $t_{0}$, the relationship $\|\bm{W}_{t}-\bm{W}_{t_{0}}\|^{2}\propto|t-t_{0}|^{c}$ holds, where $c$ may differ from $1$.
-To verify this, we measured the evolution of the squared Euclidean distance in both the weights $\bm{W}$ and the log-likelihood vectors $\bm{q}$, over steps 120k to 130k of Pythia.
-As shown in Fig.
-3, both $\bm{W}$ and $\bm{q}$ follow a power-law relationship with exponents $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ respectively.
-The weights $\bm{W}$ exhibit Brownian-like diffusion ($c_{\mathrm{w}}\approx 1$), whereas the diffusion of $\bm{q}$ is substantially suppressed ($c_{\mathrm{q}}\approx 0.2$).
-In addition, Fig.
-4 illustrates how the exponents $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ vary with the starting step.
-Notably, $c_{\mathrm{q}}$ remains consistently smaller than $c_{\mathrm{w}}$ throughout.
-Both $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ tend to decrease as training progresses.
-However, around step 120k, when the KL divergence between successive checkpoints stabilizes, $c_{\mathrm{w}}$ increases, while $c_{\mathrm{q}}$ shows no significant change.
-As discussed in Appendix B, the difference in dynamics between these two spaces could be explained by the flat minima hypothesis Keskar et al.
-(2017) and by constraints imposed by the loss function on the log-likelihood space.
-4 Fine-tuning
-Figure 5: Map showing base models and their fine-tuned variants.
-Lines represent parent-child relationships between models.
-Colors indicate model types as defined in Oyama et al.
-(2025).
-Generation refers to the depth from the root model within each lineage.
-The introduction of efficient fine-tuning methods Hu et al.
-(2022); Rafailov et al.
-(2023) has led to the development of a large number of fine-tuned models.
-In this section, we investigate the KL divergence between fine-tuned and base models.
-Settings.
-Out of the 1,018 language models analyzed in Oyama et al.
-(2025), we use 66 pairs of fine-tuned and base models (87 models in total; Fig.
-5).
-The parent-child relationships are identified using Hugging Face Hub API.
-See Appendix E for details.
-KL divergence.
-Table 1(b) shows the statistics on the KL divergence between fine-tuned and base models, and between randomly selected models.
-For the cumulative distribution, see Fig.
-18 in Appendix E.
-KL divergence values tend to be slightly smaller for models of type mistral or llama-3.
-In all cases, the KL divergence between fine-tuned and base models is smaller than that between random pairs, with a median ranging from 0.2 to 0.4 bit/byte.
-5 Layer
-Figure 6: t-SNE visualization of trajectories across layers for 50 models, obtained by applying the logit lens to each layer.
-Deeper layers are represented with lighter colors, and the final layer is indicated by a white square.
-Figure 7: For 10 models per model type: (left) t-SNE visualization of trajectories across layers, where the final layer is indicated by a white square; (right) KL divergence between adjacent layers (corresponding to Table 1(c)).
-See Fig.
-20 in Appendix F for a 3D plot.
-We investigate model trajectories across layers by treating the network up to each layer as an individual model using the logit lens nostalgebraist (2020).
-Settings.
-We use a subset of the 1,018 language models analyzed in Oyama et al.
-(2025).
-First, without restricting model types, we select the 50 most downloaded models (Fig.
-6).
-When limiting to specific model types, we use the 10 most downloaded models from each of the llama-2, llama-3, and mistral with 32 layers (30 models in total; Fig.
-7).
-See Appendix F for details on the models used.
-KL divergence.
-The right panel of Fig.
-7 and Table 1(c) show the KL divergence between consecutive layers.
-Models of the same type exhibit similar trends.
-For llama-2, Wendler et al.
-(2024) reported a sharp drop in the entropy of next-token distributions at intermediate layers in a translation task, which corresponds to the peak positions in Fig.
-7.
-The KL divergence between consecutive layers ranges from as low as 1 bit/byte to as high as 10 to 25 bit/byte, indicating substantial variation.
-Trajectory.
-Fig.
-6 and the left panel of Fig.
-7 illustrate the model trajectories across layers.
-The trajectories exhibit jumps at KL divergence peaks and tend to align in deeper layers.
-The trajectories align with increasing mean log-likelihood (Fig.
-19, Appendix F).
-6 Conclusion
-We analyzed KL divergence and trajectories across checkpoints, model pairs, and layers, revealing spiral and aligned structures over training and depth.
-Limitations
-  • The checkpoints for Pythia are available only at 1k-step intervals except for the early stages of training.
-When checkpoints are saved at shorter intervals, the resulting trajectories can differ from our observation.
-In particular, the estimation of the diffusion exponent may be affected by errors introduced by the coarse granularity of the 1k-step resolution.
-  • The KL divergence we measured depends on the text set used. In future work, we plan to investigate how the KL divergence and model trajectories vary when a different text set is used.
-  • The number of models used is limited in the experiment measuring KL divergence between fine-tuned and base models.
-This is because some fine-tuned models do not include information about their base models in their model cards obtained by using Hugging Face Hub API.
-  • In the layer-wise analysis, the logits obtained from shallow layers via the logit lens contain substantial noise.
-This issue could be addressed by using the tuned lens Belrose et al.
-(2023), an improved version of the logit lens.
-However, the number of publicly available pretrained tuned lenses is currently limited.
-We believe this limitation can be overcome in the future by training tuned lenses ourselves.
-  • While we observe structured behaviors such as spiral trajectories and subdiffusion in the log-likelihood space, their underlying causes and implications for learning dynamics have not been quantitatively analyzed and remain open for future investigation.
-Acknowledgments
-We thank Luis Iván Hernández Ruíz for helpful discussions. This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST BOOST JPMJBS2407.
-Appendix A Related Work on Oscillatory Trajectories in Weight Space
-We analyzed model trajectories in the log-likelihood space, where each model was represented by its output probabilities on a fixed text set.
-This space reflects the model’s behavior directly, and our study was, to the best of our knowledge, the first to reveal structural patterns in training dynamics within this representation.
-In particular, we observed the emergence of spiral-shaped trajectories as well as indications of constrained diffusion, which we further examine in Appendix B.
-While our main focus was on the log-likelihood space, we also briefly examined model trajectories in weight space for comparison (Appendix D.5).
-As shown in Fig.
-16, oscillatory behavior was also observed in the weight space during the later stages of training.
-Although the patterns differed in periodicity and structure, the presence of oscillations in both spaces suggests a potentially broader phenomenon in the dynamics of neural networks.
-The observed difference may reflect the presence of multiple modes of variation, with different modes being emphasized in the weight and output spaces.
-To contextualize these findings, we review prior work that has reported oscillatory behaviors in weight space during training.
-The Edge of Stability phenomenon Cohen et al.
-(2021); Arora et al.
-(2022) refers to the behavior of training loss oscillating while decreasing, observed when gradient descent enters an unstable region where the largest Hessian eigenvalue hovers near the critical learning rate.
-Singh et al.
-(2025) observed micro-level oscillations in training trajectories through changes in the cosine similarity of weights.
-Netay (2024) empirically identified double-helical trajectories in neural network weights, noting a relationship between these trajectories’ periodicity and the hyperparameters of the Adam optimizer.
-Kunin et al.
-(2024) found that weights exhibit anomalous diffusion after convergence and demonstrated oscillatory dynamics in the weight phase space, governed by the eigenstructure of the Hessian of the loss function.
-Appendix B Related Work and Discussions on Diffusion Exponents
-In Section 3, we estimated the diffusion exponents $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ corresponding to the trajectories of models in weight and log-likelihood spaces, respectively.
-Our analysis revealed that during the early phase of training, both exponents are relatively large ($c_{\mathrm{w}}=1.2$ and $c_{\mathrm{q}}=0.8$), but they decrease as training progresses.
-At the later stages, the exponent in weight space approaches $c_{\mathrm{w}}\approx 1$, which corresponds to Brownian motion, while the exponent in log-likelihood space becomes significantly smaller ($c_{\mathrm{q}}\ll 1$), which is characteristic of subdiffusion.
-In our experimental setting, texts were sampled from a corpus that closely approximates the training data.
-Consequently, each element of the log-likelihood vector represents a term contributing to the cross-entropy loss, and the negative average of these terms serves as an unbiased estimator of the expected loss.
-The contrast between the relatively large diffusion exponent in weight space and the much smaller exponent in log-likelihood space suggests that even as model weights continue to change, the output probabilities remain stable.
-As a result, the loss function is largely unaffected by these weight changes.
-This observation aligns with the hypothesis that stochastic gradient descent favors flat minima Keskar et al.
-(2017).
-Moreover, near the loss minima, the sum of negative log-likelihoods is constrained as cross-entropy loss, implying convergence to regions where many texts have high log-likelihood values.
-The observed small diffusion exponent suggests that the movement of the vectors becomes restricted in these regions.
-Appendix C Standard Error of the Estimated KL Divergence
+Geometric interpretation via folding.
+In addition to such optimization-based explanations, a more structural interpretation arises from the intrinsic redundancies of multilayer neural networks.
+Due to permutation symmetries of hidden units, many distinct weight configurations correspond to the same function or output distribution Fukumizu and Amari (2000).
+From a statistical viewpoint, such permutation-induced redundancies are associated with degenerate directions of the Fisher information matrix, giving rise to flat directions in the loss landscape.
+As a result, trajectories that are well separated in weight space can be mapped to nearby or identical points in the log-likelihood space.
+Viewed through this many-to-one mapping, training trajectories in the log-likelihood space are strongly folded, leading to complex geometric structures that remain confined to a relatively narrow region of the space.
+Quantifying folding via Hölder regularity.
+The geometric interpretation based on folding can be further quantified by considering the regularity of the mapping $f:\bm{W}\to\bm{q}$ from weight space to log-likelihood space.
+Specifically, recall that a mapping is $\alpha$-Hölder continuous if distances in the output space are bounded by a constant times the $\alpha$-th power of distances in the input space, with $\alpha=1$ corresponding to Lipschitz continuity and smaller values of $\alpha>0$ indicating increasing non-smoothness.
+If $f$ is $\alpha$-Hölder continuous and satisfies the non-degeneracy condition required to saturate the theoretical bound, then the fractal dimensions of the corresponding trajectories obey $D_{\mathrm{q}}=D_{\mathrm{w}}/\alpha$ Falconer (2014), which implies
+$\alpha=\frac{D_{\mathrm{w}}}{D_{\mathrm{q}}}=\frac{c_{\mathrm{q}}}{c_{\mathrm{w}}}.$
+This relation shows that as $\alpha$ decreases, the mapping can strongly fold the weight-space trajectory into a geometrically more complex and higher-dimensional trajectory in the log-likelihood space, while its large-scale displacement remains strongly suppressed.
+In our empirical setting, the observed diffusion exponents imply effective dimensions $D_{\mathrm{w}}\approx 2$ and $D_{\mathrm{q}}\approx 10$, which in turn yield an effective Hölder exponent of $\alpha\approx 0.2$.
+A constructive analogue for folding in log-likelihood space.
+Although the mapping $f$ induced by a neural network is explicitly defined by its architecture and parameters, its geometric regularity and folding structure are not directly accessible in closed form.
+This folding mechanism can nevertheless be illustrated by a constructive example provided by generalized Takagi–Landsberg functions.
+These functions are defined as multiscale summations of sawtooth functions $S(x)=\text{dist}(x,\mathbb{Z})$:
+$f(\bm{W})=\sum_{k=1}^{\infty}\mathbf{a}_{k}\lambda^{-k\alpha}S\!\left(\lambda^{k}\mathbf{b}_{k}^{\top}\bm{W}\right),$
+where $\lambda>1$ is a constant and $\mathbf{a}_{k}$ and $\mathbf{b}_{k}$ are vectors with the same dimensions as $\bm{q}$ and $\bm{W}$, respectively.
+This class of functions is known to be $\alpha$-Hölder continuous for any $0<\alpha<1$ under mild boundedness and non-degeneracy conditions on $\{\mathbf{a}_{k},\mathbf{b}_{k}\}$ Shidfar and Sabetfakhri (1986); Allaart and Kawamura (2011).
+Such constructions provide a simple geometric illustration of how the many-to-one and redundant parameterization of deep neural networks can give rise to highly folded trajectories in the log-likelihood space.
+In contrast to this construction, previous studies have focused on folding phenomena in input–output mappings of deep networks, where depth induces increasingly fine sawtooth partitions of the input space Telgarsky (2016, 2017).
+Appendix C On the Estimation of the KL Divergence
+Interpretation of KL divergence via text entropy.
+As a reference for interpreting KL divergence values, we estimate the entropy of the text from log-likelihood statistics, as described below.
+In our setting, this estimated text entropy is approximately 0.71 bits/byte; for example, a KL divergence of 0.1 bits/byte corresponds to a value slightly over 10% of the text entropy.
+In this work, we define the minimum value of the negative mean log-likelihood across models as an estimate of the text entropy.
+This is because, given a model $p_{i}$, a text set $\{x_{1},\ldots,x_{N}\}$, the log-likelihood $\ell_{i}(x_{s})$, and the true data-generating distribution $p_{0}$, the following approximation holds:
+$-\frac{1}{N}\sum_{s=1}^{N}\ell_{i}(x_{s})$
+$\quad\approx\mathbb{E}_{x\sim p_{0}}[-\log p_{i}(x)]$
+$\quad=\mathbb{E}_{x\sim p_{0}}[-\log p_{0}(x)]+\mathrm{KL}(p_{0},p_{i}).$
+Therefore, the minimum of the negative mean log-likelihood across models can be considered a good approximation of the text entropy.
+Standard error of the estimated KL divergence.
 Let $\bm{q}_{i}$ and $\bm{q}_{j}$ denote the $\bm{q}$ coordinates of models $p_{i}$ and $p_{j}$, respectively.
-While we use $\bm{q}$ pre-scaled by $1/\sqrt{N}$ in the main body of this paper, we do not apply this scaling in this appendix.
-Accordingly, the KL divergence between the models is estimated as $\mathrm{KL}(p_{i},p_{j})\approx\frac{1}{N}\|\bm{q}_{i}-\bm{q}_{j}\|^{2}=\frac{1}{N}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{2},$ which corresponds to the sample mean of $\{(Q_{ik}-Q_{jk})^{2}\}_{k=1}^{N}$.
+While in the main text of this paper we use $\bm{q}$ coordinates rescaled by $1/\sqrt{2N\bar{B}\log 2}$ so that squared Euclidean distances approximate KL divergence in bits per byte, we do not apply this rescaling in this appendix.
+Accordingly, the KL divergence between the models is estimated as $\mathrm{KL}(p_{i},p_{j})\approx\frac{1}{2N}\|\bm{q}_{i}-\bm{q}_{j}\|^{2}=\frac{1}{2N}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{2},$ which corresponds to the sample mean of $\{(Q_{ik}-Q_{jk})^{2}/2\}_{k=1}^{N}$.
 Assuming that the effects of double-centering in $\bm{Q}$ are negligible, these terms can be treated as approximately independent.
@@
 $\mathrm{Var}(\mathrm{KL}(p_{i},p_{j}))$
-$\approx\frac{1}{N}\left\{\frac{1}{N}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}-\frac{1}{N^{2}}\|\bm{q}_{i}-\bm{q}_{j}\|^{4}\right\}$
-$\approx\frac{1}{N^{2}}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}-\frac{1}{N}\mathrm{KL}(p_{i},p_{j})^{2}.$
+$\approx\frac{1}{N}\left\{\frac{1}{4N}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}\right$.
+$\qquad\left.-\frac{1}{4N^{2}}\|\bm{q}_{i}-\bm{q}_{j}\|^{4}\right\}$
+$\approx\frac{1}{4N^{2}}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}-\frac{1}{N}\mathrm{KL}(p_{i},p_{j})^{2}.$
 Therefore, the standard error of the estimator $\mathrm{KL}(p_{i},p_{j})$ can be approximated as
 $\mathrm{SE}(\mathrm{KL}(p_{i},p_{j}))$
-$\approx\sqrt{\frac{1}{N^{2}}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}-\frac{1}{N}\mathrm{KL}(p_{i},p_{j})^{2}}.$
-Appendix D Details of Section 3
+$\approx\sqrt{\frac{1}{4N^{2}}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}-\frac{1}{N}\mathrm{KL}(p_{i},p_{j})^{2}}.$
+The standard error estimated in nats above is converted to bits per byte by dividing by $\bar{B}\log 2$.
+Equivalently, when using $\bm{q}$ coordinates rescaled by $1/\sqrt{2N\bar{B}\log 2}$ as in the main text, the standard error of KL divergence in bits per byte is given by
+$\mathrm{SE}(\widetilde{\mathrm{KL}}(p_{i},p_{j}))$
+$\approx\sqrt{\sum_{k=1}^{N}(\tilde{Q}_{ik}-\tilde{Q}_{jk})^{4}-\frac{1}{N}\widetilde{\mathrm{KL}}(p_{i},p_{j})^{2}},$
+where $\tilde{Q}=Q/\,\sqrt{2N\bar{B}\log 2}$ and
+$\widetilde{\mathrm{KL}}(p_{i},p_{j})=\frac{\mathrm{KL}(p_{i},p_{j})}{\bar{B}\log 2}=\sum_{k=1}^{N}(\tilde{Q}_{ik}-\tilde{Q}_{jk})^{2}.$
+On the robustness to the sampling of texts.
+As an additional experiment, we computed the log-likelihood vectors using a new set of 10,000 texts randomly sampled from the 5,703,791 texts in a subset of the Pile [^6] The full Pile corpus, when segmented into 1,024-byte text chunks, yields a total of 5,703,791 texts Oyama et al.
+(2025)., using the Pythia-410M model.
+We then computed the pairwise KL divergence between all checkpoints after warmup, and found that the Pearson correlation coefficient between the KL values obtained using the original text set and those obtained using the new text set was 0.99.
+This indicates that the impact of text sampling on the KL values is limited.
+One advantage of using log-likelihood vectors is that the resolution for each domain can be adjusted by the amount of text from that domain included in the dataset.
+If a specific domain is fixed, it effectively corresponds to projecting the trajectory onto a certain aspect, and thus the shape of the trajectory is expected to change.
+Appendix D Additional Visual Analyses of Pretraining Trajectories
+This appendix provides supplementary visual analyses of pretraining trajectories.
+Appendix D.1 and Appendix D.2 focus on trajectories in the log-likelihood space, using three-dimensional t-SNE and PCA to examine their continuity and global structure, respectively.
+In contrast, Appendix D.3 presents trajectory visualizations in the weight space, highlighting fundamental limitations of joint visualization across models.
+D.1 Three-Dimensional t-SNE Visualizations
+(a) Visualization of pretraining trajectories for five Pythia model sizes using 3D t-SNE (perplexity = 30).
+This figure corresponds to a three-dimensional version of Fig.
+1, excluding additional random seeds of the 410M model.
+(b) PC2 versus PC3 for the later portion of the Pythia-2.8B training trajectory after applying 3D t-SNE.
+Figure 6: Pythia training trajectories visualized using 3D t-SNE.
+Figure 7: Autocorrelation functions of PC3 for the later portion of the Pythia-2.8B training trajectory (from step 67k onward) after applying 3D t-SNE, shown for different perplexity values.
+This subsection presents three-dimensional t-SNE visualizations of the pretraining trajectories discussed in Section 4.
+In the main text, Fig.
+1 presents two-dimensional t-SNE visualizations of pretraining trajectories for Pythia models across multiple model sizes and random seeds.
+When multiple trajectories are embedded into two dimensions, apparent discontinuities may arise near regions where trajectories overlap.
+This behavior reflects a geometric limitation of low-dimensional embeddings: in general, embedding multiple curves without crossings requires a space of dimension three or higher.
+As shown in Fig.
+6, this issue is resolved by visualizing the trajectories using three-dimensional t-SNE.
+In the 3D embedding, each training trajectory can be visualized as a continuous curve without crossings, supporting the continuity of trajectories in the original high-dimensional log-likelihood space.
+For clarity, we exclude additional seeds of the Pythia-410M model and visualize the trajectories of the five model sizes simultaneously in three dimensions, as shown in Fig.
+6.
+Figure 6 further visualizes the later portion of the Pythia-2.8B trajectory by plotting PC2 versus PC3 after applying 3D t-SNE.
+In this representation, the trajectory appears to exhibit a spiral-like structure.
+To analyze this phenomenon, Fig.
+7 shows the autocorrelation function of PC3 for the later portion of the trajectory, computed under different t-SNE perplexity settings (20, 30, 40, and 50).
+The spiral period is estimated as the training step at which the autocorrelation function attains its maximum between the second and third zero-crossings.
+As the perplexity increases, the estimated period also increases, indicating that the apparent spiral structure depends on the neighborhood size used in the embedding.
+This spiral-like appearance can be attributed to the geometric constraint of embedding high-dimensional, intricately folded trajectories into three dimensions without crossings.
+Such folding is consistent with the fractal-like structure suggested by the power-law scaling of the training trajectories discussed in the main text.
+Accordingly, the spiral pattern should be interpreted as a visualization artifact rather than as an intrinsic property of the training dynamics.
+D.2 PCA-Based Visualizations of Pretraining Trajectories
+Figure 8: PCA visualization of the training trajectories of Pythia models. (Top) PC1 vs. PC2. (Bottom) PC11 vs. PC12.
+Figure 9: Variation of principal component values from PCA applied to the trajectories of all Pythia model sizes.
+We further analyze the same pretraining trajectories using principal component analysis (PCA) to highlight their global geometric structure.
+In contrast to t-SNE, which emphasizes local neighborhood relationships, PCA captures large-scale variations shared across different model sizes.
+Fig.
+8 shows the result of applying PCA simultaneously to all model sizes and plotting their trajectories.
+Fig.
+9 further illustrates the evolution of multiple principal component values along the training steps.
+Unlike t-SNE, the top principal components exhibit similar values across model sizes, resulting in overlapping trajectories.
+This difference likely arises from the fact that the top principal components capture global variation in the trajectories, whereas t-SNE emphasizes local structure.
+In contrast, the lower principal components exhibit more noise-like behavior.
+As shown in Table 2, the KL divergence between different model sizes at the final checkpoint is larger than that between consecutive saved checkpoints.
+This is consistent with the observation that trajectories are separated in the lower components.
+D.3 Limitations of Trajectory Visualization in Weight Space
+Figure 10: t-SNE visualization (perplexity=30) of the trajectories based on pairwise distances in the weight space for Pythia models after step 100k.
+To align plot ranges across model sizes, the trajectories were centered after t-SNE.
+They were also rotated to roughly align the directions of progression.
+Final checkpoints are indicated using distinct markers.
+Since pretraining trajectories in the weight space do not admit a joint visualization across different model sizes and random seeds, we visualize them separately for each model.
+For each model size, we computed a distance matrix containing the pairwise Euclidean distances between the weight vectors of all checkpoints.
+Figure 10 shows the trajectories from step 100k onward, visualized using t-SNE777t-SNE can take a distance matrix as input..
+The trajectories were centered after t-SNE to align the plot ranges.
+In addition, for the 1B model, the trajectory was rotated by 180 degrees to roughly align its direction of progression with those of the other model sizes.
+Despite these adjustments, the resulting visualizations remain difficult to interpret.
+Because each model requires a separate embedding, the relative positions and orientations of trajectories across models are arbitrary and cannot be meaningfully compared.
+This highlights a fundamental limitation of weight-space representations for trajectory visualization, in contrast to the log-likelihood space, which provides a shared coordinate system for joint analysis.
+Appendix E Details of Sections 3.1 and 4
 To analyze the trajectories of language models during pretraining, we used Pythia models [^8] Released under the Apache 2.0 License.
-of various sizes, as listed in Table LABEL:tab:model_list_pretraining in Appendix G.
+of various sizes, as listed in Table LABEL:tab:model_list_pretraining in Appendix I.
 For each checkpoint, we computed log-likelihood vectors using the same set of 10,000 texts as in Oyama et al.
@@
 The computations were performed on an NVIDIA RTX 6000 Ada and took approximately 10 minutes for a 7B model loaded in float16 precision.
-D.1 Preprocessing
-Figure 8: Squared Euclidean distance between the weights of consecutive checkpoints for Pythia 1B. The distances between 115k and 116k, and between 116k and 117k are abnormally large.
-Figure 9: The relationship between the maximum and the standard deviation of the absolute differences in log-likelihoods between consecutive checkpoints of the same model size, computed for each text.
-Figure 10: Change in KL divergence after step 10k as a function of the number of top outlier texts removed (10, 100, 200, …, 1000). Brighter colors indicate larger numbers of removed texts.
-Assuming that the KL divergence between consecutive checkpoints changes smoothly, we regarded certain checkpoints and texts as outliers.
+E.1 Preprocessing
+Figure 11: Squared Euclidean distance between the weights of consecutive saved checkpoints for Pythia 1B. The distances between 115k and 116k, and between 116k and 117k are abnormally large.
+Figure 12: The relationship between the maximum and the standard deviation of the absolute differences in log-likelihoods between consecutive saved checkpoints of the same model size, computed for each text.
+Figure 13: Change in KL divergence after step 10k for different numbers of top outlier texts removed (10, 100, 200, …, 1000). Brighter colors indicate larger numbers of removed texts.
+Assuming that the KL divergence between consecutive saved checkpoints changes continuously, we regarded certain checkpoints and texts as outliers.
 All experiments were conducted after removing these outliers.
@@
 As shown in Fig.
-8, the Euclidean distance between weights near step 116k of Pythia 1B is abnormally large.
+11, the Euclidean distance between weights near step 116k of Pythia 1B is abnormally large.
 We suspect that this checkpoint may not have been saved properly and therefore excluded it from our analysis.
@@
 (2025), we clip the log-likelihood matrix $\bm{L}$ at the bottom 2%, after excluding the outlier model.
-Next, we consider the log-likelihood difference for each text between consecutive checkpoints after the warmup phase.
+Next, we consider the log-likelihood difference for each text between consecutive saved checkpoints after the warmup phase.
 For each text, we define its outlier score as the maximum absolute difference across all model sizes and training steps.
-Specifically, letting $p_{i,t}$ denote the model of size $i$ at step $t$, the outlier score for a text $x$ is given by $\max_{i,t}\left\{\left|p_{i,t+1}(x)-p_{i,t}(x)\right|\right\}$.
+Specifically, letting $p_{i,t}$ denote the model of size $i$ at step $t$, the outlier score for a text $x$ is given by $\max_{i,t}\left\{\left|\log p_{i,t+1}(x)-\log p_{i,t}(x)\right|\right\}$.
 It is also possible to assign scores to texts using the standard deviation of these differences.
@@
 Fig.
-10 shows the KL divergence between consecutive checkpoints after step 10k [^8] The maximum is computed from step 2k onward, after the warmup phase, but we visualize from step 10k to better capture the variation in KL divergence., plotted as a function of the number of top outlier-scored texts removed (top 10, 100, 200, …, 1000).
+13 shows the KL divergence between consecutive saved checkpoints after step 10k [^10] The maximum is computed from step 2k onward, after the warmup phase, but we visualize from step 10k to better capture the variation in KL divergence., shown for different numbers of top outlier-scored texts removed (top 10, 100, 200, …, 1000).
 When the top 300 texts (3%) are removed, the variation in KL divergence stabilizes with respect to the number of removed texts.
 Therefore, we excluded these 300 texts from our experiments.
-D.2 KL Divergence between Checkpoints
-(a) KL divergence between consecutive checkpoints during the warmup phase. The horizontal axis is scaled as $\log_{10}(1+\mathrm{Step})$.
-(b) KL divergence between consecutive checkpoints after the warmup phase. When the vertical axis is plotted on a logarithmic scale, this corresponds to the right panel of Fig. 4.
-Figure 11: KL divergence between consecutive checkpoints of Pythia. For example, the value at step 10k on the horizontal axis represents the KL divergence between the checkpoints at steps 10k and 11k.
-Figure 12: KL divergence and its standard error between consecutive checkpoints of Pythia after step 10k.
+E.2 KL Divergence between Checkpoints
+Figure 14: KL divergence and its standard error between consecutive saved checkpoints of Pythia after step 10k.
 410M 1B 1.4B 2.8B 6.9B
@@
 6.9B 1.7 0.92 0.57 0.21 0.00
-Table 3: KL divergence (bit/byte) between different model sizes at the final checkpoint of Pythia.
-Fig.
-11 shows the variation in KL divergence between consecutive checkpoints for all model sizes.
-In this figure, the value at step 10k, for example, represents the KL divergence between the checkpoints at steps 10k and 11k.
+Table 2: KL divergence (bits/byte) between different model sizes at the final checkpoint of Pythia.
 During pretraining, the learning rate changes dynamically with the training step.
 In particular, the first 1,430 steps correspond to the warmup phase, during which the learning rate increases linearly.
-As a result, as shown in Fig.
-11, the KL divergence between the checkpoints takes on relatively large values, mostly exceeding 10 bit/byte during this interval.
+As a result, as shown in the left panel of Fig.
+4, the KL divergence between the checkpoints takes on relatively large values, mostly exceeding 10 bits/byte during this interval.
+The right panel of Fig.
+4 shows the variation in KL divergence between consecutive saved checkpoints for all model sizes.
+In this figure, the value at step 10k, for example, represents the KL divergence between the checkpoints at steps 10k and 11k.
 Additionally, Fig.
-12 plots the KL divergence between consecutive checkpoints after step 10k for Pythia, along with the corresponding standard error.
+14 plots the KL divergence between consecutive saved checkpoints after step 10k for Pythia, along with the corresponding standard error.
 The derivation of the standard error is described in Appendix C.
@@
 As discussed in Section 4, the divergence caused by differences in model size is larger than that resulting from 1k steps of training.
-D.3 Period of Spiral Trajectory
-Figure 13: (Left) PC2 vs. PC3 after applying PCA to the later half portion of the 3D t-SNE trajectories for each Pythia model size. (Right) Autocorrelation function of PC3.
-In this appendix, we describe the procedure used to analyze the spiral trajectories suggested in Fig.
-1.
-For each model size in Pythia, 3D t-SNE is applied independently.
-PCA is then performed on the sequence of checkpoints from the midpoint of the trajectory (step 67k onward) to extract principal axes.
-Since PC1 corresponds to the overall direction of trajectory progression, we plot PC2 versus PC3 in the left panel of Fig.
-13, where the trajectory approximately forms a circular pattern.
-We also compute the autocorrelation function of the PC3 values, treated as a time series, and show the result in the right panel of Fig.
-13.
-The spiral period reported in Table 2 is defined as the step at which the autocorrelation reaches its maximum between the second and third zero-crossings.
-D.4 PCA
-Figure 14: PCA visualization of the training trajectories of Pythia models. (Top) PC1 vs. PC2. (Bottom) PC11 vs. PC12.
-Figure 15: Variation of principal component values from PCA applied to the trajectories of all Pythia model sizes.
-In Section 3, we visualized the trajectories of Pythia during pretraining using t-SNE. In this appendix, we instead visualize the trajectories using PCA.
-Fig.
-14 shows the result of applying PCA simultaneously to all model sizes and plotting their trajectories.
-Fig.
-15 further illustrates the evolution of multiple principal component values along the training steps.
-Unlike t-SNE, the top principal components exhibit similar values across model sizes, resulting in overlapping trajectories.
-This difference likely arises from the fact that the top principal components capture global variation in the trajectories, whereas t-SNE emphasizes local structure.
-In contrast, the lower principal components exhibit more noise-like behavior.
-As shown in Table 3, the KL divergence between different model sizes at the final checkpoint is larger than that between consecutive checkpoints.
-This is consistent with the observation that trajectories are separated in the lower components.
-D.5 Comparison with Weights
-Figure 16: t-SNE visualization of the trajectories based on weight-space distances for Pythia models after step 100k.
-To align scales across model sizes, the trajectories were centered after t-SNE.
-They were also rotated to align the directions of progression.
-Final checkpoints are indicated using distinct markers.
+E.3 Estimation of Diffusion Exponents
 Size Log-likelihood Weight
@@
 Table 3: The mean and standard deviation of the coefficient of determination ($R^{2}$) when varying the starting step used to compute the diffusion exponent.
-This appendix describes the procedure for comparing language models using the Euclidean distance between the weights of Pythia checkpoints.
-Trajectory.
-For each model size, we computed a distance matrix containing the pairwise Euclidean distances among all checkpoints.
-Fig.
-16 shows the trajectories from step 100k onward, visualized using t-SNE999t-SNE can take a distance matrix as input..
-The trajectories were centered after t-SNE to align the plot ranges.
-In addition, for the 1B model, the trajectory was rotated by 180 degrees to align its rough direction of progression with that of the other model sizes.
-Anomalous diffusion.
-Fig.
+In this subsection, we describe the procedure for estimating diffusion exponents in both the log-likelihood and weight spaces.
+The left panel of Fig.
 5 shows the evolution of squared Euclidean distances in weights and log-likelihood vectors over 10k steps starting from step 120k.
-Fig.
+The right panel of Fig.
 5 illustrates how the estimated diffusion exponent changes when varying the starting step for the calculation.
@@
 Summary statistics of the coefficient of determination ($R^{2}$) for each regression are provided in Table 3.
-D.6 Different Seeds
+E.4 Different Seeds
 Figure 15: KL divergence between the final checkpoints of Pythia 410M models with different seeds.
-van der Wal et al.
-(2025) trained Pythia models with nine different seeds and released checkpoints at the same steps as in the original Pythia suite.
+wal2025polypythias trained Pythia models with nine different seeds and released checkpoints at the same steps as in the original Pythia suite.
 The random seed affects the initialization of weights and the order of the training data.
@@
 Despite sharing the same architecture and data, the KL divergence between seed 3 and the other seeds, as well as between seed 4 and the others, are significantly larger.
-According to van der Wal et al.
-(2025), models trained with seed 3 and seed 4 experienced loss spikes during training and are considered outliers in terms of performance.
+According to wal2025polypythias, models trained with seed 3 and seed 4 experienced loss spikes during training and are considered outliers in terms of performance.
 This suggests that KL divergence is also effective in identifying such outlier models.
@@
 1.
-We also report in Table 1(a) the average KL divergence between the final checkpoint of the original 410M model and those of seeds 1, 2, 5, 6, 7, 8, and 9.
-Appendix E Details of Section 4
-Figure 18: Cumulative distribution of KL divergence for pairs of fine-tuned models and their base models, and for randomly selected model pairs (corresponding to Table 1(b)).
-The horizontal axis is scaled as $\log_{10}(1+\mathrm{KL})$.
-The dotted lines indicate the median of each distribution.
-We describe the language models used in Section 4.
+We also report in Table 1(b) the average KL divergence between the final checkpoint of the original 410M model and those of seeds 1, 2, 5, 6, 7, 8, and 9.
+Appendix F Details of Section 3.2
+Here, we describe the language models used in Section 3.2.
+We selected the 50 most downloaded models from those analyzed in Oyama et al.
+(2025).
+A complete list of the models is provided in Table LABEL:tab:model_list_layer in Appendix I.
+For the quantization methods, we applied 8-bit and 4-bit quantization using bitsandbytes [^12] https://github.com/bitsandbytes-foundation/bitsandbytes Dettmers et al. (2022).
+Appendix G Details of Section 3.3
+We describe the language models used in Section 3.3.
 From the 1,018 models analyzed in Oyama et al.
@@
 Models associated with multiple base models (i.e., merged models) were excluded.
-A list of the language models used in this analysis is provided in Table LABEL:tab:model_list_ft in Appendix G.
+A list of the language models used in this analysis is provided in Table LABEL:tab:model_list_ft in Appendix I.
 In Fig.
-5, there are a few pairs connected by line segments that span across different colors, namely different model types.
+2b, there are a few pairs connected by line segments that span across different colors, namely different model types.
 This occurs because, under the model type classification method used in Oyama et al.
 (2025), some fine-tuned models were categorized as belonging to the “others” type.
-In our experiments, when we refer to, for example, a “llama-3 fine-tuned pair”, we mean a pair in which the fine-tuned model is of the llama-3 type.
+Summary statistics for these distributions are reported in Table 1(a).
 Fig.
-18 illustrates the cumulative distributions of KL divergence for pairs of fine-tuned models and their base models, as well as for randomly selected model pairs.
-Summary statistics for these distributions are reported in Table 1(b).
-The figure shows that KL divergence after fine-tuning tends to be slightly smaller, as indicated by the left-skewed distribution.
+3 shows that KL divergence after fine-tuning tends to be slightly smaller, as indicated by the left-skewed distribution.
 As noted in Oyama et al.
@@
 To more accurately assess the effect of fine-tuning, we also analyze random model pairs sampled within the same model type.
-Appendix F Details of Section 5
-Figure 19: Re-rendering of the top panel of Fig. 6. Colors indicate mean log-likelihood, clipped at the bottom 5% across all values to improve visibility.
-Figure 20: 3D t-SNE visualization of the left panel of Fig. 7. Deeper layers are represented with lighter colors, and the final layer is indicated by a white square.
-Here, we describe the language models used in Section 5.
+Appendix H Details of Section 3.4
+Figure 16: Re-rendering of the top panel of Fig. 2c. Colors indicate mean log-likelihood, clipped at the bottom 5% across all values to improve visibility.
+Figure 17: For 10 models per model type: (left) t-SNE visualization of trajectories across layers, where the final layer is indicated by a white square; (right) KL divergence between adjacent layers (corresponding to Table 1(c)).
+Figure 18: 3D t-SNE visualization (perplexity=30) of the left panel of Fig. 17. Deeper layers are represented with lighter colors, and the final layer is indicated by a white square.
+Here, we describe the language models used in Section 3.4.
 For the trajectories shown in Fig.
-6 and Fig.
-19, we selected the 50 most downloaded models from those analyzed in Oyama et al.
+2c and Fig.
+16, we selected the 50 most downloaded models from those analyzed in Oyama et al.
 (2025).
-For specified model type analysis, we selected models with 32 layers from the llama-2, llama-3, and mistral types.
-For each of these types, we used the 10 most downloaded models.
-These models were used in the analysis shown in Fig.
-7 and Fig.
-20.
-A complete list of the models is provided in Table LABEL:tab:model_list_layer in Appendix G.
+A complete list of the models is provided in Table LABEL:tab:model_list_layer in Appendix I.
 Fig.
@@
 This indicates that the direction of trajectory progression corresponds to deeper layers, that is, to increasing mean log-likelihood.
-Additionally, Fig.
-20 presents a 3D version of the left panel of Fig.
-6.
-This illustration implies that the jump observed in Fig.
-7 is not an artifact of the 2D t-SNE visualization, but rather results from a genuinely large KL divergence between adjacent layers.
-Appendix G Model List
-The models used in Sections 3, 4, and 5 are listed in Tables LABEL:tab:model_list_pretraining, LABEL:tab:model_list_ft, and LABEL:tab:model_list_layer, respectively.
+Additionally, for detailed analysis, we selected models with 32 layers from the llama-2, llama-3, and mistral types.
+For each of these types, we used the 10 most downloaded models.
+That is, the top 50 do not include 10 models for each of llama-2, llama-3, and mistral types.
+Therefore, additional models were included to ensure 10 models per category, resulting in a total of 9 added models.
+Fig.
+17 shows the trajectories of these models and the KL divergence between adjacent layers, while Fig.
+18 illustrates the trajectories in 3d space.
+For all model types, the trajectories are already clustered by model type from the shallow layers onward.
+In addition, the KL divergence values exhibit similar trends within each type.
+Appendix I Model List
+The models used in Sections 3.1, 3.3, and 3.4 are listed in Tables LABEL:tab:model_list_pretraining, LABEL:tab:model_list_ft, and LABEL:tab:model_list_layer, respectively.
 Table LABEL:tab:model_list_pretraining is sorted by model size; Table LABEL:tab:model_list_ft is sorted alphabetically by model name, considering the model generation, model type, and corresponding base model; and Table LABEL:tab:model_list_layer is sorted in descending order of download counts.
@@
 (2025).
-Note that, for the Pythia models trained with different random seeds van der Wal et al.
-(2025) and used in Section 3, the corresponding BibTeX entries in Table LABEL:tab:model_list_pretraining were prepared manually.
-Table 5: List of 14 models in the experiments for pretraining models. “ID” denotes the index sorted by model size; “Model Name” denotes the name of the model.
+Note that, for the Pythia models trained with different random seeds wal2025polypythias and used in Section 4, the corresponding BibTeX entries in Table LABEL:tab:model_list_pretraining were prepared manually.
+Table 4: List of 14 models in the experiments for pretraining models.
+“ID” denotes the index sorted by model size; “Model Name” denotes the name of the model.
+Seeds 3 and 4 of Pythia-410M are excluded from the experiments in Sections 3.1 and 4.
 ID Model Name
-1 EleutherAI/pythia-410m Gao et al. (2020); Biderman et al. (2022, 2023)
-2 EleutherAI/pythia-410m-seed1 van der Wal et al. (2025)
-3 EleutherAI/pythia-410m-seed2 van der Wal et al. (2025)
-4 EleutherAI/pythia-410m-seed3 van der Wal et al. (2025)
-5 EleutherAI/pythia-410m-seed4 van der Wal et al. (2025)
-6 EleutherAI/pythia-410m-seed5 van der Wal et al. (2025)
-7 EleutherAI/pythia-410m-seed6 van der Wal et al. (2025)
-8 EleutherAI/pythia-410m-seed7 van der Wal et al. (2025)
-9 EleutherAI/pythia-410m-seed8 van der Wal et al. (2025)
-10 EleutherAI/pythia-410m-seed9 van der Wal et al. (2025)
-11 EleutherAI/pythia-1b Gao et al. (2020); Biderman et al. (2022, 2023)
-12 EleutherAI/pythia-1.4b Gao et al. (2020); Biderman et al. (2022, 2023)
-13 EleutherAI/pythia-2.8b Gao et al. (2020); Biderman et al. (2022, 2023)
-14 EleutherAI/pythia-6.9b Gao et al. (2020); Biderman et al. (2022, 2023)
+1 EleutherAI/pythia-410m arxiv:2101.00027; arxiv:2201.07311; arxiv:2304.01373
+2 EleutherAI/pythia-410m-seed1 wal2025polypythias
+3 EleutherAI/pythia-410m-seed2 wal2025polypythias
+4 EleutherAI/pythia-410m-seed3 wal2025polypythias
+5 EleutherAI/pythia-410m-seed4 wal2025polypythias
+6 EleutherAI/pythia-410m-seed5 wal2025polypythias
+7 EleutherAI/pythia-410m-seed6 wal2025polypythias
+8 EleutherAI/pythia-410m-seed7 wal2025polypythias
+9 EleutherAI/pythia-410m-seed8 wal2025polypythias
+10 EleutherAI/pythia-410m-seed9 wal2025polypythias
+11 EleutherAI/pythia-1b arxiv:2101.00027; arxiv:2201.07311; arxiv:2304.01373
+12 EleutherAI/pythia-1.4b arxiv:2101.00027; arxiv:2201.07311; arxiv:2304.01373
+13 EleutherAI/pythia-2.8b arxiv:2101.00027; arxiv:2201.07311; arxiv:2304.01373
+14 EleutherAI/pythia-6.9b arxiv:2101.00027; arxiv:2201.07311; arxiv:2304.01373
 Table 5: List of 87 models in the experiments for fine-tuning models.
@@
 ID Model Name Base Model Name Model Type Generation DLs
-1 deepseek-ai/DeepSeek-Prover-V1.5-Base Xin et al. (2024) deepseek 1 230
-2 tiiuae/falcon-rw-1b Brown et al. (2020); Press et al. (2022); Dao et al. (2022); Penedo et al. (2023) falcon 1 22921
-3 google/gemma-1.1-7b-it Joshi et al.
-(2017); Mihaylov et al.
-(2018); Zhao et al.
-(2018); Clark et al.
-(2019); Talmor et al.
-(2019); Sap et al.
-(2019); Sakaguchi et al.
-(2019); Bisk et al.
-(2019); Chollet (2019); Zellers et al.
-(2019); Hendrycks et al.
-(2021); Chen et al.
-(2021); Austin et al.
-(2021); Cobbe et al.
-(2021); Parrish et al.
-(2022); Srivastava et al.
-(2023); Zhong et al.
-(2023); Gemini Team et al.
-(2024) gemma 1 19835
-4 google/gemma-7b Joshi et al.
-(2017); Mihaylov et al.
-(2018); Zhao et al.
-(2018); Rudinger et al.
-(2018); Bisk et al.
-(2019); Chollet (2019); Clark et al.
-(2019); Talmor et al.
-(2019); Sap et al.
-(2019); Zellers et al.
-(2019); Sakaguchi et al.
-(2019); Gehman et al.
-(2020); Chen et al.
-(2021); Austin et al.
-(2021); Hendrycks et al.
-(2021); Dhamala et al.
-(2021); Cobbe et al.
-(2021); Lin et al.
-(2022); Hartvigsen et al.
-(2022); Parrish et al.
-(2022); Srivastava et al.
-(2023); Zhong et al.
-(2023); Dettmers et al.
-(2023); Gemini Team et al.
-(2024) gemma 1 72047
-5 openchat/openchat-3.5-0106-gemma Wang et al. (2024a) gemma 1 7817
-6 openlm-research/open_llama_3b Together Computer (2023); Touvron et al. (2023a); Geng and Liu (2023) llama-1 1 157405
-7 codellama/CodeLlama-13b-Instruct-hf Rozière et al. (2024) llama-2 1 25248
-8 meta-llama/Llama-2-13b-hf Touvron et al. (2023b) llama-2 1 120871
-9 meta-llama/Llama-2-7b-hf Touvron et al. (2023b) llama-2 1 1294737
-10 meta-llama/Meta-Llama-3-8B-Instruct AI@Meta (2024) llama-3 1 2058011
-11 meta-llama/Meta-Llama-3-8B AI@Meta (2024) llama-3 1 675850
-12 NousResearch/Hermes-2-Pro-Llama-3-8B "Teknium" et al. (2024b) llama-3 1 26797
-13 mistralai/Mistral-7B-Instruct-v0.2 Jiang et al. (2023) mistral 1 3565248
-14 mistralai/Mistral-7B-v0.1 Jiang et al. (2023) mistral 1 1750089
+1 deepseek-ai/DeepSeek-Prover-V1.5-Base arxiv:2408.08152 deepseek 1 230
+2 tiiuae/falcon-rw-1b arxiv:2005.14165; arxiv:2108.12409; arxiv:2205.14135; arxiv:2306.01116 falcon 1 22921
+3 google/gemma-1.1-7b-it arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1907.10641; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.07830; arxiv:2009.03300; arxiv:2107.03374; arxiv:2108.07732; arxiv:2110.14168; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2312.11805 gemma 1 19835
+4 google/gemma-7b arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1804.09301; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1905.07830; arxiv:1907.10641; arxiv:2009.11462; arxiv:2107.03374; arxiv:2108.07732; arxiv:2009.03300; arxiv:2101.11718; arxiv:2110.14168; arxiv:2109.07958; arxiv:2203.09509; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2305.14314; arxiv:2312.11805 gemma 1 72047
+5 openchat/openchat-3.5-0106-gemma arxiv:2309.11235 gemma 1 7817
+6 openlm-research/open_llama_3b together2023redpajama; arxiv:2302.13971; openlm2023openllama llama-1 1 157405
+7 codellama/CodeLlama-13b-Instruct-hf arxiv:2308.12950 llama-2 1 25248
+8 meta-llama/Llama-2-13b-hf arxiv:2307.09288 llama-2 1 120871
+9 meta-llama/Llama-2-7b-hf arxiv:2307.09288 llama-2 1 1294737
+10 meta-llama/Meta-Llama-3-8B-Instruct llama3modelcard llama-3 1 2058011
+11 meta-llama/Meta-Llama-3-8B llama3modelcard llama-3 1 675850
+12 NousResearch/Hermes-2-Pro-Llama-3-8B Hermes-2-Pro-Llama-3-8B llama-3 1 26797
+13 mistralai/Mistral-7B-Instruct-v0.2 arxiv:2310.06825 mistral 1 3565248
+14 mistralai/Mistral-7B-v0.1 arxiv:2310.06825 mistral 1 1750089
 15 mlabonne/Marcoro14-7B-slerp mistral 1 3792
 16 mlabonne/NeuralMonarch-7B mistral 1 13486
-17 stabilityai/japanese-stablelm-base-gamma-7b Jiang et al. (2023) mistral 1 2072
+17 stabilityai/japanese-stablelm-base-gamma-7b arxiv:2310.06825 mistral 1 2072
 18 beomi/KoRWKV-6B others 1 2127
-19 microsoft/Orca-2-13b Mitra et al. (2023) others 1 13616
-20 upstage/SOLAR-10.7B-v1.0 Kim et al. (2024b) others 1 24478
-21 01-ai/Yi-1.5-9B 01. AI et al. (2025) others 1 20252
-22 deepseek-ai/DeepSeek-Prover-V1.5-SFT Xin et al. (2024) deepseek-ai/DeepSeek-Prover-V1.5-Base deepseek 2 6459
+19 microsoft/Orca-2-13b arxiv:2311.11045 others 1 13616
+20 upstage/SOLAR-10.7B-v1.0 arxiv:2312.15166 others 1 24478
+21 01-ai/Yi-1.5-9B arxiv:2403.04652 others 1 20252
+22 deepseek-ai/DeepSeek-Prover-V1.5-SFT arxiv:2408.08152 deepseek-ai/DeepSeek-Prover-V1.5-Base deepseek 2 6459
 23 euclaise/falcon_1b_stage1 tiiuae/falcon-rw-1b falcon 2 2119
@@
 25 lemon-mint/gemma-7b-openhermes-v0.80 google/gemma-1.1-7b-it gemma 2 4867
-26 google/gemma-7b-it Joshi et al.
-(2017); Mihaylov et al.
-(2018); Zhao et al.
-(2018); Rudinger et al.
-(2018); Bisk et al.
-(2019); Chollet (2019); Clark et al.
-(2019); Talmor et al.
-(2019); Sap et al.
-(2019); Zellers et al.
-(2019); Sakaguchi et al.
-(2019); Gehman et al.
-(2020); Chen et al.
-(2021); Austin et al.
-(2021); Hendrycks et al.
-(2021); Dhamala et al.
-(2021); Cobbe et al.
-(2021); Lin et al.
-(2022); Hartvigsen et al.
-(2022); Parrish et al.
-(2022); Srivastava et al.
-(2023); Zhong et al.
-(2023); Gemini Team et al.
-(2024) google/gemma-7b gemma 2 66776
+26 google/gemma-7b-it arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1804.09301; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1905.07830; arxiv:1907.10641; arxiv:2009.11462; arxiv:2107.03374; arxiv:2108.07732; arxiv:2009.03300; arxiv:2101.11718; arxiv:2110.14168; arxiv:2109.07958; arxiv:2203.09509; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2312.11805 google/gemma-7b gemma 2 66776
 27 Telugu-LLM-Labs/Indic-gemma-7b-finetuned-sft-Navarasa-2.0 google/gemma-7b gemma 2 2025
 28 lemon-mint/gemma-ko-7b-instruct-v0.62 openchat/openchat-3.5-0106-gemma gemma 2 7543
-29 rinna/youri-7b Andonian et al. (2021); Touvron et al. (2023b); Zhao et al. (2023); Sawada et al. (2024) meta-llama/Llama-2-7b-hf llama-2 2 2512
-30 aaditya/Llama3-OpenBioLLM-8B Singhal et al. (2022); Nori et al. (2023); Singhal et al. (2023); Pal and Sankarasubbu (2024b, a); Rafailov et al. (2024) meta-llama/Meta-Llama-3-8B llama-3 2 9308
+29 rinna/youri-7b gpt-neox-library; arxiv:2307.09288; rinna-youri-7b; arxiv:2404.01657 meta-llama/Llama-2-7b-hf llama-2 2 2512
+30 aaditya/Llama3-OpenBioLLM-8B arxiv:2212.13138; arxiv:2303.13375; arxiv:2305.09617; OpenBioLLMs; arxiv:2402.07023; arxiv:2305.18290 meta-llama/Meta-Llama-3-8B llama-3 2 9308
 31 cognitivecomputations/dolphin-2.9-llama3-8b meta-llama/Meta-Llama-3-8B llama-3 2 130977
@@
 36 jondurbin/bagel-8b-v1.0 meta-llama/Meta-Llama-3-8B llama-3 2 7960
-37 openchat/openchat-3.6-8b-20240522 Wang et al. (2024a) meta-llama/Meta-Llama-3-8B llama-3 2 10464
-38 rinna/llama-3-youko-8b Andonian et al. (2021); Mitsuda et al. (2024); AI@Meta (2024); Sawada et al. (2024) meta-llama/Meta-Llama-3-8B llama-3 2 1468
+37 openchat/openchat-3.6-8b-20240522 arxiv:2309.11235 meta-llama/Meta-Llama-3-8B llama-3 2 10464
+38 rinna/llama-3-youko-8b gpt-neox-library; rinna-llama-3-youko-8b; llama3modelcard; arxiv:2404.01657 meta-llama/Meta-Llama-3-8B llama-3 2 1468
 39 ruslanmv/Medical-Llama3-8B meta-llama/Meta-Llama-3-8B llama-3 2 5457
-40 lightblue/suzume-llama-3-8B-multilingual Devine (2024) meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 16442
+40 lightblue/suzume-llama-3-8B-multilingual arxiv:2405.12612 meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 16442
 41 MaziyarPanahi/Llama-3-8B-Instruct-v0.4 meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 1407
-42 PathFinderKR/Waktaverse-Llama-3-KO-8B-Instruct AI@Waktaverse (2024); AI@Meta (2024) meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 2305
-43 shenzhi-wang/Llama3-8B-Chinese-Chat Wang et al. (2024b) meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 55718
+42 PathFinderKR/Waktaverse-Llama-3-KO-8B-Instruct waktaversellama3modelcard; llama3modelcard meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 2305
+43 shenzhi-wang/Llama3-8B-Chinese-Chat shenzhi-wang-2024 meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 55718
 44 SJ-Donald/llama3-passthrough-chat meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 2241
-45 swap-uniba/LLaMAntino-3-ANITA-8B-Inst-DPO-ITA Basile et al. (2023); AI@Meta (2024); Polignano et al. (2024) meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 5995
-46 NousResearch/Hermes-2-Theta-Llama-3-8B "Teknium" et al. (2024a) NousResearch/Hermes-2-Pro-Llama-3-8B llama-3 2 12392
-47 vicgalle/Configurable-Hermes-2-Pro-Llama-3-8B Gallego (2024) NousResearch/Hermes-2-Pro-Llama-3-8B llama-3 2 10273
+45 swap-uniba/LLaMAntino-3-ANITA-8B-Inst-DPO-ITA arxiv:2312.09993; llama3modelcard; arxiv:2405.07101 meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 5995
+46 NousResearch/Hermes-2-Theta-Llama-3-8B Hermes-2-Theta-Llama-3-8B NousResearch/Hermes-2-Pro-Llama-3-8B llama-3 2 12392
+47 vicgalle/Configurable-Hermes-2-Pro-Llama-3-8B arxiv:2404.00495 NousResearch/Hermes-2-Pro-Llama-3-8B llama-3 2 10273
 48 abacusai/bigstral-12b-32k mistralai/Mistral-7B-Instruct-v0.2 mistral 2 5216
@@
 51 cognitivecomputations/dolphin-2.2.1-mistral-7b mistralai/Mistral-7B-v0.1 mistral 2 7292
-52 cypienai/cymist-2-v02-SFT Lacoste et al. (2019) mistralai/Mistral-7B-v0.1 mistral 2 2717
+52 cypienai/cymist-2-v02-SFT arxiv:1910.09700 mistralai/Mistral-7B-v0.1 mistral 2 2717
 53 HuggingFaceH4/mistral-7b-sft-beta mistralai/Mistral-7B-v0.1 mistral 2 7408
-54 HuggingFaceH4/zephyr-7b-alpha Ding et al. (2023); Tunstall et al. (2023); Cui et al. (2024); Rafailov et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 12911
-55 HuggingFaceH4/zephyr-7b-beta Ding et al. (2023); Tunstall et al. (2023); Cui et al. (2024); Rafailov et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 294019
-56 ibm/merlinite-7b Sudalairaj et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 11156
+54 HuggingFaceH4/zephyr-7b-alpha arxiv:2305.14233; arxiv:2310.16944; arxiv:2310.01377; arxiv:2305.18290 mistralai/Mistral-7B-v0.1 mistral 2 12911
+55 HuggingFaceH4/zephyr-7b-beta arxiv:2305.14233; arxiv:2310.16944; arxiv:2310.01377; arxiv:2305.18290 mistralai/Mistral-7B-v0.1 mistral 2 294019
+56 ibm/merlinite-7b arxiv:2403.01081 mistralai/Mistral-7B-v0.1 mistral 2 11156
 57 INSAIT-Institute/BgGPT-7B-Instruct-v0.2 mistralai/Mistral-7B-v0.1 mistral 2 3265
-58 Intel/neural-chat-7b-v3-1 Mukherjee et al. (2023) mistralai/Mistral-7B-v0.1 mistral 2 3976
-59 kaist-ai/mistral-orpo-capybara-7k Hong et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 4821
+58 Intel/neural-chat-7b-v3-1 arxiv:2306.02707 mistralai/Mistral-7B-v0.1 mistral 2 3976
+59 kaist-ai/mistral-orpo-capybara-7k arxiv:2403.07691 mistralai/Mistral-7B-v0.1 mistral 2 4821
 60 Locutusque/Hercules-2.5-Mistral-7B mistralai/Mistral-7B-v0.1 mistral 2 1953
-61 mistralai/Mistral-7B-Instruct-v0.1 Jiang et al. (2023) mistralai/Mistral-7B-v0.1 mistral 2 1370245
-62 NousResearch/Hermes-2-Pro-Mistral-7B "interstellarninja" et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 14586
-63 NousResearch/Nous-Hermes-2-Mistral-7B-DPO "Teknium" et al. (2024c) mistralai/Mistral-7B-v0.1 mistral 2 8725
-64 openchat/openchat-3.5-0106 Wang et al. (2024a); OpenAI et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 26858
-65 openchat/openchat-3.5-1210 Wang et al. (2024a); OpenAI et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 2123
+61 mistralai/Mistral-7B-Instruct-v0.1 arxiv:2310.06825 mistralai/Mistral-7B-v0.1 mistral 2 1370245
+62 NousResearch/Hermes-2-Pro-Mistral-7B Hermes-2-Pro-Mistral-7B mistralai/Mistral-7B-v0.1 mistral 2 14586
+63 NousResearch/Nous-Hermes-2-Mistral-7B-DPO Nous-Hermes-2-Mistral-7B-DPO mistralai/Mistral-7B-v0.1 mistral 2 8725
+64 openchat/openchat-3.5-0106 arxiv:2309.11235; arxiv:2303.08774 mistralai/Mistral-7B-v0.1 mistral 2 26858
+65 openchat/openchat-3.5-1210 arxiv:2309.11235; arxiv:2303.08774 mistralai/Mistral-7B-v0.1 mistral 2 2123
 66 statking/zephyr-7b-sft-full-orpo mistralai/Mistral-7B-v0.1 mistral 2 2278
@@
 71 beomi/KoAlpaca-KoRWKV-6B beomi/KoRWKV-6B others 2 2288
-72 Nexusflow/NexusRaven-V2-13B Nexusflow.ai team (2023); Rozière et al. (2024) codellama/CodeLlama-13b-Instruct-hf others 2 3966
-73 haoranxu/ALMA-13B-Pretrain Xu et al. (2024a, b) meta-llama/Llama-2-13b-hf others 2 5510
+72 Nexusflow/NexusRaven-V2-13B nexusraven; arxiv:2308.12950 codellama/CodeLlama-13b-Instruct-hf others 2 3966
+73 haoranxu/ALMA-13B-Pretrain arxiv:2309.11674; arxiv:2401.08417 meta-llama/Llama-2-13b-hf others 2 5510
 74 ruslanmv/ai-medical-model-32bit meta-llama/Meta-Llama-3-8B-Instruct others 2 2810
@@
 79 NousResearch/Nous-Hermes-2-SOLAR-10.7B upstage/SOLAR-10.7B-v1.0 others 2 9918
-80 upstage/SOLAR-10.7B-Instruct-v1.0 Kim et al. (2024b, a) upstage/SOLAR-10.7B-v1.0 others 2 67725
+80 upstage/SOLAR-10.7B-Instruct-v1.0 arxiv:2312.15166; arxiv:2403.19270 upstage/SOLAR-10.7B-v1.0 others 2 67725
 81 cognitivecomputations/dolphin-2.9.1-yi-1.5-9b 01-ai/Yi-1.5-9B others 2 4880
-82 deepseek-ai/DeepSeek-Prover-V1.5-RL Xin et al. (2024) deepseek-ai/DeepSeek-Prover-V1.5-SFT deepseek 3 12223
+82 deepseek-ai/DeepSeek-Prover-V1.5-RL arxiv:2408.08152 deepseek-ai/DeepSeek-Prover-V1.5-SFT deepseek 3 12223
 83 euclaise/falcon_1b_stage2 euclaise/falcon_1b_stage1 falcon 3 4016
@@
 85 argilla/notus-7b-v1 alignment-handbook/zephyr-7b-sft-full mistral 3 7660
-86 Intel/neural-chat-7b-v3-3 Yu et al. (2024) Intel/neural-chat-7b-v3-1 mistral 3 166226
+86 Intel/neural-chat-7b-v3-3 arxiv:2309.12284 Intel/neural-chat-7b-v3-1 mistral 3 166226
 87 MaziyarPanahi/Llama-3-8B-Instruct-v0.9 MaziyarPanahi/Llama-3-8B-Instruct-v0.8 llama-3 4 6241
-Table 7: List of 59 models in the experiments for layer-wise models.
+Table 6: List of 50 (+9) models in the experiments for quantization and layer-wise models.
 “ID” denotes the index sorted in descending order of downloads; “Model Name” denotes the name of the model; “Model Type” denotes the model classification defined in Oyama et al.
@@
 ID Model Name Model Type Layers DLs
-1 mistralai/Mistral-7B-Instruct-v0.2 Jiang et al. (2023) mistral 32 3565248
-2 meta-llama/Meta-Llama-3-8B-Instruct AI@Meta (2024) llama-3 32 2058011
-3 mistralai/Mistral-7B-v0.1 Jiang et al. (2023) mistral 32 1750089
+1 mistralai/Mistral-7B-Instruct-v0.2 arxiv:2310.06825 mistral 32 3565248
+2 meta-llama/Meta-Llama-3-8B-Instruct llama3modelcard llama-3 32 2058011
+3 mistralai/Mistral-7B-v0.1 arxiv:2310.06825 mistral 32 1750089
 4 mistralai/Mistral-7B-v0.3 mistral 32 1443065
-5 meta-llama/Llama-2-7b-chat-hf Touvron et al. (2023b) llama-2 32 1402244
-6 mistralai/Mistral-7B-Instruct-v0.1 Jiang et al. (2023) mistral 32 1370245
-7 meta-llama/Llama-2-7b-hf Touvron et al. (2023b) llama-2 32 1294737
+5 meta-llama/Llama-2-7b-chat-hf arxiv:2307.09288 llama-2 32 1402244
+6 mistralai/Mistral-7B-Instruct-v0.1 arxiv:2310.06825 mistral 32 1370245
+7 meta-llama/Llama-2-7b-hf arxiv:2307.09288 llama-2 32 1294737
 8 TinyLlama/TinyLlama-1.1B-Chat-v1.0 others 22 1078500
 9 TinyLlama/TinyLlama-1.1B-intermediate-step-1431k-3T others 22 798206
-10 meta-llama/Meta-Llama-3-8B AI@Meta (2024) llama-3 32 675850
+10 meta-llama/Meta-Llama-3-8B llama3modelcard llama-3 32 675850
 11 OpenAssistant/oasst-sft-4-pythia-12b-epoch-3.5 gpt_neox 36 458718
-12 bigcode/starcoder2-3b Beltagy et al. (2020); Dao et al. (2022); Bavarian et al. (2022); Ainslie et al. (2023); Lozhkov et al. (2024) others 30 445316
-13 lmsys/vicuna-7b-v1.5 Touvron et al. (2023b); Zheng et al. (2023) llama-2 32 368410
-14 HuggingFaceH4/zephyr-7b-beta Ding et al. (2023); Tunstall et al. (2023); Cui et al. (2024); Rafailov et al. (2024) mistral 32 294019
-15 meta-llama/Llama-2-13b-chat-hf Touvron et al. (2023b) llama-2 40 266090
-16 google/gemma-2b Joshi et al.
-(2017); Mihaylov et al.
-(2018); Zhao et al.
-(2018); Rudinger et al.
-(2018); Bisk et al.
-(2019); Chollet (2019); Clark et al.
-(2019); Talmor et al.
-(2019); Sap et al.
-(2019); Zellers et al.
-(2019); Sakaguchi et al.
-(2019); Gehman et al.
-(2020); Chen et al.
-(2021); Austin et al.
-(2021); Hendrycks et al.
-(2021); Dhamala et al.
-(2021); Cobbe et al.
-(2021); Lin et al.
-(2022); Hartvigsen et al.
-(2022); Parrish et al.
-(2022); Srivastava et al.
-(2023); Zhong et al.
-(2023); Gemini Team et al.
-(2024) gemma 18 256798
-17 EleutherAI/gpt-neo-1.3B Gao et al. (2020); Black et al. (2021) others 24 243332
-18 EleutherAI/gpt-j-6b Gao et al. (2020); Wang and Komatsuzaki (2021); Wang (2021); Su et al. (2023) gptj 28 241435
+12 bigcode/starcoder2-3b arxiv:2004.05150; arxiv:2205.14135; arxiv:2207.14255; arxiv:2305.13245; arxiv:2402.19173 others 30 445316
+13 lmsys/vicuna-7b-v1.5 arxiv:2307.09288; arxiv:2306.05685 llama-2 32 368410
+14 HuggingFaceH4/zephyr-7b-beta arxiv:2305.14233; arxiv:2310.16944; arxiv:2310.01377; arxiv:2305.18290 mistral 32 294019
+15 meta-llama/Llama-2-13b-chat-hf arxiv:2307.09288 llama-2 40 266090
+16 google/gemma-2b arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1804.09301; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1905.07830; arxiv:1907.10641; arxiv:2009.11462; arxiv:2107.03374; arxiv:2108.07732; arxiv:2009.03300; arxiv:2101.11718; arxiv:2110.14168; arxiv:2109.07958; arxiv:2203.09509; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2312.11805 gemma 18 256798
+17 EleutherAI/gpt-neo-1.3B arxiv:2101.00027; gpt-neo others 24 243332
+18 EleutherAI/gpt-j-6b arxiv:2101.00027; gpt-j; mesh-transformer-jax; arxiv:2104.09864 gptj 28 241435
 19 microsoft/phi-2 others 32 237268
-20 EleutherAI/gpt-neo-2.7B Gao et al. (2020); Black et al. (2021) others 32 205503
+20 EleutherAI/gpt-neo-2.7B arxiv:2101.00027; gpt-neo others 32 205503
 21 huggyllama/llama-7b llama-1 32 192383
-22 tiiuae/falcon-7b-instruct Shazeer (2019); Brown et al. (2020); Dao et al. (2022); Su et al. (2023); Almazrouei et al. (2023); Penedo et al. (2023) falcon 32 179952
+22 tiiuae/falcon-7b-instruct arxiv:1911.02150; arxiv:2005.14165; arxiv:2205.14135; arxiv:2104.09864; falcon40b; arxiv:2306.01116 falcon 32 179952
 23 DeepMount00/Llama-3-8b-Ita llama-3 32 179401
-24 Intel/neural-chat-7b-v3-3 Yu et al. (2024) mistral 32 166226
-25 openlm-research/open_llama_3b Together Computer (2023); Touvron et al. (2023a); Geng and Liu (2023) llama-1 26 157405
+24 Intel/neural-chat-7b-v3-3 arxiv:2309.12284 mistral 32 166226
+25 openlm-research/open_llama_3b together2023redpajama; arxiv:2302.13971; openlm2023openllama llama-1 26 157405
 26 augmxnt/shisa-gamma-7b-v1 mistral 32 151426
-27 facebook/opt-1.3b Brown et al. (2020); Zhang et al. (2022) opt 24 139758
-28 Qwen/Qwen1.5-1.8B Bai et al. (2023) others 24 137256
-29 codellama/CodeLlama-7b-Instruct-hf Rozière et al. (2024) llama-2 32 133523
+27 facebook/opt-1.3b arxiv:2005.14165; arxiv:2205.01068 opt 24 139758
+28 Qwen/Qwen1.5-1.8B arxiv:2309.16609 others 24 137256
+29 codellama/CodeLlama-7b-Instruct-hf arxiv:2308.12950 llama-2 32 133523
 30 cognitivecomputations/dolphin-2.9-llama3-8b llama-3 32 130977
-31 Qwen/Qwen1.5-7B Bai et al. (2023) others 32 122768
-32 meta-llama/Llama-2-13b-hf Touvron et al. (2023b) llama-2 40 120871
-33 microsoft/phi-1_5 Li et al. (2023) others 24 112476
-34 NousResearch/Meta-Llama-3-8B-Instruct AI@Meta (2024) llama-3 32 104388
-35 tiiuae/falcon-7b Shazeer (2019); Brown et al. (2020); Gao et al. (2020); Dao et al. (2022); Su et al. (2023); Almazrouei et al. (2023); Penedo et al. (2023) falcon 32 104010
-36 google/gemma-2b-it Joshi et al.
-(2017); Mihaylov et al.
-(2018); Zhao et al.
-(2018); Rudinger et al.
-(2018); Bisk et al.
-(2019); Chollet (2019); Clark et al.
-(2019); Talmor et al.
-(2019); Sap et al.
-(2019); Zellers et al.
-(2019); Sakaguchi et al.
-(2019); Gehman et al.
-(2020); Chen et al.
-(2021); Austin et al.
-(2021); Hendrycks et al.
-(2021); Dhamala et al.
-(2021); Cobbe et al.
-(2021); Lin et al.
-(2022); Hartvigsen et al.
-(2022); Parrish et al.
-(2022); Srivastava et al.
-(2023); Zhong et al.
-(2023); Gemini Team et al.
-(2024) gemma 18 103851
+31 Qwen/Qwen1.5-7B arxiv:2309.16609 others 32 122768
+32 meta-llama/Llama-2-13b-hf arxiv:2307.09288 llama-2 40 120871
+33 microsoft/phi-1_5 arxiv:2309.05463 others 24 112476
+34 NousResearch/Meta-Llama-3-8B-Instruct llama3modelcard llama-3 32 104388
+35 tiiuae/falcon-7b arxiv:1911.02150; arxiv:2005.14165; arxiv:2101.00027; arxiv:2205.14135; arxiv:2104.09864; falcon40b; arxiv:2306.01116 falcon 32 104010
+36 google/gemma-2b-it arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1804.09301; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1905.07830; arxiv:1907.10641; arxiv:2009.11462; arxiv:2107.03374; arxiv:2108.07732; arxiv:2009.03300; arxiv:2101.11718; arxiv:2110.14168; arxiv:2109.07958; arxiv:2203.09509; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2312.11805 gemma 18 103851
 37 teknium/OpenHermes-2.5-Mistral-7B mistral 32 100997
-38 mosaicml/mpt-7b-chat Henry et al. (2020); Press et al. (2022); Dao et al. (2022); MosaicML NLP Team (2023) others 32 88069
-39 Qwen/CodeQwen1.5-7B-Chat Bai et al. (2023) others 32 77169
-40 GritLM/GritLM-7B Muennighoff et al. (2025) mistral 32 72991
-41 google/gemma-7b Joshi et al.
-(2017); Mihaylov et al.
-(2018); Zhao et al.
-(2018); Rudinger et al.
-(2018); Bisk et al.
-(2019); Chollet (2019); Clark et al.
-(2019); Talmor et al.
-(2019); Sap et al.
-(2019); Zellers et al.
-(2019); Sakaguchi et al.
-(2019); Gehman et al.
-(2020); Chen et al.
-(2021); Austin et al.
-(2021); Hendrycks et al.
-(2021); Dhamala et al.
-(2021); Cobbe et al.
-(2021); Lin et al.
-(2022); Hartvigsen et al.
-(2022); Parrish et al.
-(2022); Srivastava et al.
-(2023); Zhong et al.
-(2023); Dettmers et al.
-(2023); Gemini Team et al.
-(2024) gemma 28 72047
-42 upstage/SOLAR-10.7B-Instruct-v1.0 Kim et al. (2024b, a) others 48 67725
-43 google/gemma-7b-it Joshi et al.
-(2017); Mihaylov et al.
-(2018); Zhao et al.
-(2018); Rudinger et al.
-(2018); Bisk et al.
-(2019); Chollet (2019); Clark et al.
-(2019); Talmor et al.
-(2019); Sap et al.
-(2019); Zellers et al.
-(2019); Sakaguchi et al.
-(2019); Gehman et al.
-(2020); Chen et al.
-(2021); Austin et al.
-(2021); Hendrycks et al.
-(2021); Dhamala et al.
-(2021); Cobbe et al.
-(2021); Lin et al.
-(2022); Hartvigsen et al.
-(2022); Parrish et al.
-(2022); Srivastava et al.
-(2023); Zhong et al.
-(2023); Gemini Team et al.
-(2024) gemma 28 66776
-44 codellama/CodeLlama-7b-hf Rozière et al. (2024) llama-2 32 66040
-45 facebook/opt-2.7b Brown et al. (2020); Zhang et al. (2022) opt 32 61450
-46 shenzhi-wang/Llama3-8B-Chinese-Chat Wang et al. (2024b) llama-3 32 55718
+38 mosaicml/mpt-7b-chat arxiv:2010.04245; arxiv:2108.12409; arxiv:2205.14135; MosaicML2023Introducing others 32 88069
+39 Qwen/CodeQwen1.5-7B-Chat arxiv:2309.16609 others 32 77169
+40 GritLM/GritLM-7B arxiv:2402.09906 mistral 32 72991
+41 google/gemma-7b arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1804.09301; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1905.07830; arxiv:1907.10641; arxiv:2009.11462; arxiv:2107.03374; arxiv:2108.07732; arxiv:2009.03300; arxiv:2101.11718; arxiv:2110.14168; arxiv:2109.07958; arxiv:2203.09509; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2305.14314; arxiv:2312.11805 gemma 28 72047
+42 upstage/SOLAR-10.7B-Instruct-v1.0 arxiv:2312.15166; arxiv:2403.19270 others 48 67725
+43 google/gemma-7b-it arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1804.09301; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1905.07830; arxiv:1907.10641; arxiv:2009.11462; arxiv:2107.03374; arxiv:2108.07732; arxiv:2009.03300; arxiv:2101.11718; arxiv:2110.14168; arxiv:2109.07958; arxiv:2203.09509; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2312.11805 gemma 28 66776
+44 codellama/CodeLlama-7b-hf arxiv:2308.12950 llama-2 32 66040
+45 facebook/opt-2.7b arxiv:2005.14165; arxiv:2205.01068 opt 32 61450
+46 shenzhi-wang/Llama3-8B-Chinese-Chat shenzhi-wang-2024 llama-3 32 55718
 47 VAGOsolutions/Llama-3-SauerkrautLM-8b-Instruct llama-3 32 55400
-48 lmsys/vicuna-13b-v1.5 Touvron et al. (2023b); Zheng et al. (2023) llama-2 40 48653
-49 Qwen/Qwen2-1.5B Yang et al. (2024) others 28 46510
-50 openlm-research/open_llama_7b Together Computer (2023); Touvron et al. (2023a); Geng and Liu (2023) llama-1 32 44968
+48 lmsys/vicuna-13b-v1.5 arxiv:2307.09288; arxiv:2306.05685 llama-2 40 48653
+49 Qwen/Qwen2-1.5B qwen2 others 28 46510
+50 openlm-research/open_llama_7b together2023redpajama; arxiv:2302.13971; openlm2023openllama llama-1 32 44968
 51 LinkSoul/Chinese-Llama-2-7b llama-2 32 40799
 52 mistral-community/Mistral-7B-v0.2 mistral 32 30074
-53 NousResearch/Hermes-2-Pro-Llama-3-8B "Teknium" et al. (2024b) llama-3 32 26797
-54 lightblue/suzume-llama-3-8B-multilingual Devine (2024) llama-3 32 16442
-55 abacusai/Llama-3-Smaug-8B Pal et al. (2024) llama-3 32 12873
+53 NousResearch/Hermes-2-Pro-Llama-3-8B Hermes-2-Pro-Llama-3-8B llama-3 32 26797
+54 lightblue/suzume-llama-3-8B-multilingual arxiv:2405.12612 llama-3 32 16442
+55 abacusai/Llama-3-Smaug-8B arxiv:2402.13228 llama-3 32 12873
 56 NousResearch/Nous-Hermes-llama-2-7b llama-2 32 12019
 57 togethercomputer/LLaMA-2-7B-32K llama-2 32 8884
-58 elyza/ELYZA-japanese-Llama-2-7b-instruct Touvron et al. (2023b); Sasaki et al. (2023) llama-2 32 6296
-59 togethercomputer/Llama-2-7B-32K-Instruct Liu et al. (2023) llama-2 32 5596
+58 elyza/ELYZA-japanese-Llama-2-7b-instruct arxiv:2307.09288; elyzallama2023 llama-2 32 6296
+59 togethercomputer/Llama-2-7B-32K-Instruct arxiv:2307.03172 llama-2 32 5596